### v1

# Hessian Spectrum at the Ground State in Spherical Spin Glasses

Abstract
We determine the bulk Hessian spectrum at approximate ground states in any spherical spin glass. Marginal stability, the presence of many near-zero Hessian eigenvalues, is shown to be equivalent to full replica symmetry breaking near overlap $1$ at zero temperature. This connection has long been implicit in the physics literature, which also links marginal stability to the performance of efficient algorithms. For even models, we prove the Hessian has no outlier eigenvalues, which has geometric consequences for low temperature Gibbs measures in the case that marginal stability is absent. Our proofs rely on interpolation bounds for certain vector spin glass models. For generic spin glasses, we give another more conceptual argument that full RSB near overlap $1$ implies marginal stability at low temperature.
1 Introduction
Spherical spin glass Hamiltonians are random disordered smooth functions in very high dimension. Their landscapes are understood to be rich and complicated, with exponentially many local maxima at a range of energy levels. We will be interested in the qualitative behavior around their near-global maxima. Does local uniform concavity hold near the extreme values, or could the Hessian be ill-conditioned? The former would imply that low temperature Gibbs measures are supported on isolated wells, and that the corresponding stationary dynamics remain trapped within them. By contrast an ill-conditioned marginally stable Hessian would allow for the possibility of a connected “manifold” of near maxima. In the physics literature, marginal stability at low-temperature is widely believed to be equivalent to full replica symmetry breaking (full RSB) near overlap $1$ at zero temperature, which is a property of the order parameter in the Parisi formula. We prove this equivalence for spherical spin glasses with even interactions, and derive consequences for Langevin dynamics and disorder chaos at low temperature whenever full RSB is absent.
As our results depend on properties of the minimizer in Parisi’s variational formula, we begin by recalling this formula. For each $p\geq 1$, let $\mathbf{G}^{(p)}\in\left({\mathbb{R}}^{N}\right)^{\otimes p}$ be an independent $p$-tensor with i.i.d. standard Gaussian entries. Fixing an infinite sequence $(\gamma_{p})_{p\geq 1}$ of non-negative reals, the mixed $p$-spin Hamiltonian $H_{N}$ is
$H_{N}({\boldsymbol{\sigma}})$ $=\sum_{p\geq 1}\frac{\gamma_{p}}{N^{(p-1)/2}}\langle\mathbf{G}^{(p)},{\boldsymbol{\sigma}}^{\otimes p}\rangle.$ (1.1)
The coefficients $\gamma_{p}$ are encoded in the mixture function $\xi(x)=\sum_{p\geq 1}\gamma_{p}^{2}x^{p}$, which we assume is not linear and has radius of convergence strictly larger than $1$. We view $H_{N}$ as a function on the spherical domain ${\mathcal{S}}_{N}\equiv\big{\{}{\boldsymbol{\sigma}}\in{\mathbb{R}}^{N}:\sum_{i=1}^{N}{\boldsymbol{\sigma}}_{i}^{2}=N\big{\}}$; equivalently, it is the centered Gaussian process on ${\mathcal{S}}_{N}$ with covariance
$\operatorname*{{\mathbb{E}}}H_{N}({\boldsymbol{\sigma}}^{1})H_{N}({\boldsymbol{\sigma}}^{2})=N\xi(R({\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2}))\equiv N\xi(\langle{\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2}\rangle/N).$ (1.2)
Here $R({\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2})=\langle{\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2}\rangle/N\in[-1,1]$ is known as the overlap between ${\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2}$.
We will study $H_{N}$ near its extreme values, where the ground state energy
${\mathrm{GS}}_{N}=\max_{{\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}}H_{N}({\boldsymbol{\sigma}})/N.$ (1.3)
is approximately achieved. The limiting value of ${\mathrm{GS}}_{N}$ is given by Parisi’s formula. Following [CS17] (see also [CS92, Tal06a, Che13]), let ${\mathcal{N}}$ denote the set of non-decreasing, right-continuous functions $\zeta:[0,1)\to{\mathbb{R}}_{\geq 0}$ and let
${\mathcal{K}}=\left\{(\zeta,L)\in{\mathcal{N}}\times(0,\infty)\leavevmode\nobreak\ :\leavevmode\nobreak\ L>\int_{0}^{1}\zeta(s)\leavevmode\nobreak\ {\rm d}s\right\}.$
For $(\zeta,L)\in{\mathcal{K}}$, define $\hat{\zeta}(q)=L-\int_{0}^{q}\zeta(s)\leavevmode\nobreak\ {\rm d}s$ and:
${\mathcal{Q}}(\zeta,L)=\frac{1}{2}\left(\xi^{\prime}(0)L+\int_{0}^{1}\xi^{\prime\prime}(q)\hat{\zeta}(q)\leavevmode\nobreak\ {\rm d}q+\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}(q)}\right)=\frac{1}{2}\left(\xi^{\prime}(1)L+\int_{0}^{1}\xi^{\prime\prime}(q)\Big{(}\int_{0}^{q}\zeta(s){\rm d}s\Big{)}\leavevmode\nobreak\ {\rm d}q+\int_{0}^{1}\frac{{\rm d}q}{\hat{\zeta}(q)}\right).$
Proposition 1.1 ([CS17]).
The in-probability limit of the ground state energy is:
${\mathrm{GS}}(\xi)\equiv\operatorname*{p-lim}_{N\to\infty}{\mathrm{GS}}_{N}=\inf_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L).$ (1.4)
Further, there exists a unique minimizing pair $(\zeta,L)\in{\mathcal{K}}$.
The minimizing $(\zeta,L)$ is characterized by local stationarity conditions, which are reviewed in Subsection 1.6.
1.1 Main Results
As previously mentioned, we study the local behavior of $H_{N}$ around its extreme values. To start, Proposition 1.2 shows the bulk Hessian spectrum at the ground state is directly described by the minimizer $(\zeta,L)\in{\mathcal{K}}$ in the zero-temperature Parisi formula (1.4), in particular the value $\hat{\zeta}(1)=L-\int_{0}^{1}\zeta(s){\rm d}s$. Namely the rescaled radial derivative and bulk spectral edges of the Hessian are asymptotically given by the formulas:
$r(\xi)$ $=\hat{\zeta}(1)\xi^{\prime\prime}(1)+\hat{\zeta}(1)^{-1},$ (1.5)
$\lambda_{\pm}(\xi)$ =±$=\pm 2\sqrt{\xi^{\prime\prime}(1)}-r(\xi)=-\hat{\zeta}(1)\left(\sqrt{\xi^{\prime\prime}(1)}\mp\hat{\zeta}(1)^{-1}\right)^{2}\leq 0.$ (1.6)
The same description extends, up to error $o_{\delta\to 0}(1)$, to all $\delta$-approximate ground states, i.e. points in the set
$A_{\delta}=\big{\{}{\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}\leavevmode\nobreak\ :\leavevmode\nobreak\ H_{N}({\boldsymbol{\sigma}})/N\geq{\mathrm{GS}}(\xi)-\delta\big{\}}$ (1.7)
Below and throughout, we write ${\boldsymbol{\lambda}}_{k}(\cdot)$ for the $k$-th largest eigenvalue of a symmetric matrix. $\nabla_{\mathrm{sp}}^{2}H_{N}(\cdot)$ denotes the Riemannian Hessian on ${\mathcal{S}}_{N}$, an $(N-1)\times(N-1)$ matrix defined in Subsection 1.5. We often say an event depending on $H_{N}$ and other parameters (e.g. $(\xi,\varepsilon,\delta)$) holds with probability (at least) $1-e^{-cN}$. Here $c$ is always sufficiently small depending on the other parameters, while $N$ is sufficiently large depending on everything else including $c$.
Proposition 1.2.
For any $\xi$ and $\varepsilon>0$, there exists $\delta>0$ such that the following holds with probability $1-e^{-cN}$ for large $N$. For all $\delta$-approximate ground states ${\boldsymbol{\sigma}}\in A_{\delta}$, the gradient satisfies
$\|\nabla H_{N}({\boldsymbol{\sigma}})-r(\xi)\,{\boldsymbol{\sigma}}\|/\sqrt{N}\leq\varepsilon$ (1.8)
while the top and bottom of the bulk Hessian spectrum satisfy
$\Big{|}{\boldsymbol{\lambda}}_{\lfloor\delta N\rfloor}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-\lambda_{+}(\xi)\Big{|}$ $\leq\varepsilon,$ (1.9)
$\Big{|}{\boldsymbol{\lambda}}_{N-\lfloor\delta N\rfloor}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-\lambda_{-}(\xi)\Big{|}$ $\leq\varepsilon.$
This result is stated as a proposition because a different formula for $\nabla H_{N}({\boldsymbol{\sigma}})$ follows using the ideas in [Sub24, Corollary 7] at zero temperature; see also [CPS23, CPS21] for Ising spin glasses. Our first step in proving (1.8) uses exactly this method, which yields a more abstract formula for $r(\xi)$ as the derivative of ${\mathrm{GS}}(\xi)$ along a perturbation of $\xi$. Applying the stationarity conditions for $(\zeta,L)$, we then deduce the more tractable (1.8). Regarding the bulk spectrum (1.9), known concentration estimates imply that uniformly for all ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, the spectral measure of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ is approximately a semi-circle density with radius $2\sqrt{\xi^{\prime\prime}(1)}$ and shifted by the rescaled radial derivative $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})=R({\boldsymbol{\sigma}},\nabla H_{N}({\boldsymbol{\sigma}}))$; see the proof of [Sub21, Lemma 3] (which is rephrased below as Lemma 1.11). Thus (1.9) follows routinely from (1.8).
Note that $\lambda_{+}$ defined above is never positive, and equals 00 if and only if
$\hat{\zeta}(1)=\xi^{\prime\prime}(1)^{-1/2}.$ (1.10)
We say $\xi$ exhibits full RSB endpoint behavior when this condition holds. As explained in Corollary 1.14, this condition is implied by (and is presumably generically equivalent to) the more standard definition of full RSB that $\zeta$ is strictly increasing in a neighborhood of $1$. When (1.10) holds, Proposition 1.2 states that the Hessian at any approximate ground state has many near-zero eigenvalues, i.e is marginally stable. Conversely when $\lambda_{+}<0$, it is natural to hope that $H_{N}$ is locally uniformly concave at near-extrema. However it is not obvious why the Hessian does not have outlier eigenvalues. Our next result proves exactly this when $\xi$ contains only even degree terms.111 Theorem 1.3 and Corollary 1.4 extend with no changes when $\gamma_{1}>0$, but do require $\gamma_{p}=0$ for odd $p\geq 3$. The same holds for Corollaries 1.5 and 1.6. However deducing slow mixing from Corollary 1.6 requires modification if $\gamma_{p}>0$, as does Corollary 1.7. Indeed for the former, one must exclude the topologically trivial phase where low-temperature Langevin dynamics does mix rapidly [HS23c, Theorem 1.8]. As discussed further in Subsection 1.4, the idea is to connect extreme Hessian eigenvalues to the ground state energy of certain vector spin glasses, which can be determined to first-order using precise interpolation bounds.
Theorem 1.3.
Suppose $\xi$ is even, i.e. $\gamma_{p}=0$ for all odd $p$. Then for any $\varepsilon$ there is $\delta$ such that the following holds with probability $1-e^{-cN}$ for large $N$. For all ${\boldsymbol{\sigma}}\in A_{\delta}$,
$\Big{|}{\boldsymbol{\lambda}}_{1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-\lambda_{+}(\xi)\Big{|}$ $\leq\varepsilon,$ (1.11)
$\Big{|}{\boldsymbol{\lambda}}_{N-1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-\lambda_{-}(\xi)\Big{|}$ $\leq\varepsilon,$ (1.12)
The following corollary is essentially immediate. (For the easy upper bound ${\boldsymbol{\lambda}}_{1}\leq\varepsilon/2$ in (1.13) without assuming $\xi$ is even, see Corollary 1.10.)
Corollary 1.4.
If $\xi$ exhibits full RSB endpoint behavior (1.10), then all approximate ground states are marginally stable. Namely for any $\varepsilon>0$, if $\delta=\delta(\xi,\varepsilon)$ is small enough, then with probability $1-e^{-cN}$ all ${\boldsymbol{\sigma}}\in A_{\delta}$ satisfy:
$|{\boldsymbol{\lambda}}_{\lfloor\delta N\rfloor}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}|+|{\boldsymbol{\lambda}}_{1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}|\leq\varepsilon.$ (1.13)
Conversely, suppose $\xi$ does not exhibit full RSB endpoint behavior and is even, and let ${\boldsymbol{\sigma}}\in A_{\delta}$ for small $\delta\leq\delta_{*}(\xi)$. Then $H_{N}$ is locally uniformly concave near ${\boldsymbol{\sigma}}$ in that with probability $1-e^{-cN}$:
${\boldsymbol{\lambda}}_{1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}\leq-1/C(\xi)<0.$ (1.14)
We note that these results apply to Gibbs samples at temperature tending to 00 slowly with $N$, simply because they are approximate ground states (see e.g. Lemma 1.12). They also have consequences at positive temperatures not tending to 00. Here as usual, the Gibbs measure $\mu_{\beta}=\mu_{\beta,H_{N}}$ at inverse temperature $\beta$ is defined by
${\rm d}\mu_{\beta}({\boldsymbol{\sigma}})=e^{\beta H_{N}({\boldsymbol{\sigma}})}{\rm d}\mu_{0}({\boldsymbol{\sigma}})/Z_{N,\beta}.$
$Z_{N,\beta}=\int_{{\mathcal{S}}_{N}}e^{\beta H_{N}({\boldsymbol{\sigma}})}{\rm d}\mu_{0}({\boldsymbol{\sigma}})$ is the partition function relative to uniform measure $\mu_{0}$ on ${\mathcal{S}}_{N}$. Under genericity conditions on $\xi$, $\mu_{\beta}$ is arranged into an ultrametric tree of ancestor states, which are certain points in the interior of ${\mathcal{S}}_{N}$ that are approximate ground states for their respective radius $t\sqrt{N}$; see [Pan13a, Jag17, CS21, Sub24]. Then applying the above results to $H_{N}|_{t{\mathcal{S}}_{N}}$, which due to the formula (1.2) amounts to studying $\xi_{t}(q)=\xi(t^{2}q)$, describes the local behavior of $H_{N}$ near ancestor states. The corresponding zero-temperature order parameter $(\zeta_{t},L_{t})$ can be directly read off from the positive-temperature analog of the original model, and will exhibit full RSB endpoint behavior if and only if the positive-temperature Parisi order parameter for $\xi$ exhibits analogous full RSB behavior at overlap $t^{2}$ (see [Sub24, Proposition 11]).
Additionally, the local uniform concavity in (1.14) easily implies a qualitative description of the deep level sets $A_{\delta}$. As stated in Corollary 1.5 below, all connected components $A_{\delta}$ consist of small separated clusters; this is a more geometric formulation of non-marginal stability. A similar result is deduced in [BSZ20, Proposition 9.1] for a special class of $\xi$, for which a version of (1.14) follows from Kac–Rice asymptotics. (The proof in our setting is strictly easier, since the input result in [BSZ20] gives Hessian control only at critical points.)
Corollary 1.5.
Suppose $\xi$ does not exhibit full RSB endpoint behavior and is even. For some $C(\xi)>0$ and all small enough $\delta\in(0,\delta_{*}(\xi))$, let $A_{\delta}$ be as in (1.7). Then with probability $1-e^{-cN}$, every connected component of $A_{\delta}$ contains exactly $1$ critical point of $H_{N}$, which is a local maximum, and has diameter at most $\sqrt{N\delta}/C(\xi)$. Further, distinct components of $A_{\delta}$ are separated by $\sqrt{N}/C(\xi)$.
We pause here to emphasize two points. First, Corollary 1.5 relies crucially on the absence of Hessian outliers stated in (1.11): even a single near 00 outlier eigenvalue could result in $A_{\delta}$ having large components. Second, the full RSB assumption is essentially necessary: when $\lambda_{+}=0$, Corollary 1.5 evidently does not hold at any ${\boldsymbol{\sigma}}\in A_{\delta}$. Indeed, it is predicted in e.g. [CK94, Section 9] that under full RSB behavior, $\delta=O(1/N)$ may suffice for $A_{\delta}$ to have components of macroscopic diameter. Rigorously, it follows (e.g. by generic perturbations as in [Pan13b, Chapter 3.7]) that when $\xi$ is full RSB near overlap $1$ in the sense that $\zeta$ from (1.4) is strictly increasing on $[q,1]$, then for any $\delta>0$ and sufficiently large $N$, with high probability there exist ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in A_{\delta}$ with $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=q$.
Corollary 1.5 yields further consequences for low temperature Gibbs measures in the absence of marginal stability. Using the separation of distinct components, and that low temperature Gibbs samples are approximate ground states, one obtains a plateau property for the autocorrelation function of stationary Langevin dynamics at low temperature. Given an initialization ${\boldsymbol{x}}_{0}\in{\mathcal{S}}_{N}$, the spherical Langevin dynamics are defined by
${\rm d}{\boldsymbol{x}}_{t}=\left(\beta\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{x}}_{t})-\frac{(N-1){\boldsymbol{x}}_{t}}{N}\right){\rm d}t+P_{{\boldsymbol{x}}_{t}}^{\perp}\sqrt{2}\,{\rm d}\boldsymbol{B}_{t}$ (1.15)
where $P_{{\boldsymbol{x}}_{t}}^{\perp}=I_{N}-{\boldsymbol{x}}_{t}^{\otimes 2}/N$ is a rank $1$ projection and $\boldsymbol{B}_{t}$ is standard $N$-dimensional Browian motion. It is well known that these dynamics remain on ${\mathcal{S}}_{N}$ almost surely, with unique stationary distribution $\mu_{\beta}$. We omit the proof of the corollary below, which is exactly the same as [AMS23a, Corollary 2.6]. A byproduct is exponentially slow mixing of Langevin dynamics under the same conditions (since $\mu_{\beta}$ is origin-symmetric when $\xi$ is even).
Corollary 1.6.
Suppose $\xi$ does not exhibit full RSB endpoint behavior and is even, and fix $\varepsilon>0$. For some $C(\xi)>0$ and all large enough $\beta\geq\beta(\xi,\varepsilon)$, let ${\boldsymbol{x}}_{0}\sim\mu_{\beta}$ be a Gibbs sample, and ${\boldsymbol{x}}_{t}$ be the trajectory of spherical Langevin dynamics (1.15). Then with probability $1-e^{-cN}$, one has that
$\inf_{0\leq t\leq e^{cN}}R({\boldsymbol{x}}_{0},{\boldsymbol{x}}_{t})\geq 1-\varepsilon.$
The last assertion of Corollary 1.5 also allows us to determine the scale for the onset of transport disorder chaos, i.e. the sensitivity of $\mu_{\beta}$ to perturbations of $H_{N}$. Here we also assume $\xi$ is “even generic”, with $\sum_{p\geq 1}\frac{1_{\gamma_{p}\neq 0}}{p}=\infty$. With $\tilde{H}_{N}$ an independent copy of $H_{N}=H_{N,0}$, set
$H_{N,t}=\sqrt{1-t}H_{N}+\sqrt{t}\tilde{H}_{N},\quad\forall\,t\in[0,1].$
Let $\mu_{\beta,t}=\mu_{\beta,H_{N,t}}$ be the corresponding Gibbs measure, and ${\mathbb{W}}_{2,N}(\mu,\mu^{\prime})$ for the rescaled Wasserstein-$2$ distance
$W_{2,N}(\mu,\mu^{\prime})^{2}=\inf_{\pi\in\Pi(\mu,\mu^{\prime})}{\mathbb{E}}^{\pi}[\|{\boldsymbol{x}}-{\boldsymbol{y}}\|_{2}^{2}/N].$
Here the infimum is over couplings $({\boldsymbol{x}},{\boldsymbol{y}})\sim\pi$ with marginals ${\boldsymbol{x}}\sim\mu$ and ${\boldsymbol{y}}\sim\mu^{\prime}$.
Corollary 1.7.
Suppose $\xi$ is even generic and does not exhibit full RSB endpoint behavior. Let $\beta\geq\beta(\xi)$ be sufficiently large. If $(\varepsilon_{N})_{N\geq 1}$ is a deterministic sequence with $\liminf_{N\to\infty}N\varepsilon_{N}>0$, then
$\liminf_{N\to\infty}{\mathbb{E}}[{\mathbb{W}}_{2,N}(\mu_{\beta,0},\mu_{\beta,\varepsilon_{N}})]>0.$ (1.16)
The same $1/N$ scaling of $\varepsilon_{N}$ was identified for pure spherical spin glasses at very low temperature [Sub17b], and established later within the shattered phase by [AMS23a]. Indeed as we explain in Subsection 2.4, Corollary 1.7 follows by the arguments in [AMS23a, Section 5.1]. We note the decay condition on $\varepsilon_{N}$ is sharp for any $(\xi,\beta)$: [AMS23a, Proposition 2.8] shows if $\varepsilon_{N}\leq o(1/N)$, then the expected total variation distance between $\mu_{\beta,0}$ and $\mu_{\beta,\varepsilon_{N}}$ tends to 00; thus (1.16) does not hold, i.e. there is no chaos. We refer to [AMS23a, Section 2.3] for further discussion, and comparison with the overlap-based notion of disorder chaos appearing in e.g. [Cha09, CS17, CHL18, Eld20].
1.2 Connections to the Kac–Rice Formula
The Kac–Rice formula [Kac43, Ric44, AT07] forms the basis for a substantial line of work on extreme values in spherical spin glasses stemming from [ABČ13, AB13]. The formula gives general expressions for the expectations (or higher moments) of the number of critical points of $H_{N}$ at different energy levels (or with other characteristics), which can be studied using random matrix theory. Notably [Sub17a, BSZ20] identified the ground state energy ${\mathrm{GS}}(\xi)$ for a subclass of “pure-like” $\xi$, by matching the first and second moments of the number of critical points at energy ${\mathrm{GS}}_{N}$. These models exhibit $1$-RSB behavior, i.e. the minimizing function $\zeta$ in the Parisi formula is a positive constant. Recently the author and Huang extended this argument to all $1$-RSB models in [HS23a], by truncating the second moment based on interpolation bounds.
Whenever the Kac–Rice formula suffices to identify the ground state, the radial derivative (and hence bulk Hessian spectrum) at approximate ground states can be read off from the calculation, and the Hessians will have no outliers in the sense of Theorem 1.3. The reason is that the expected number of local maxima at the ground state energy level without the correct radial derivative (or with outlier eigenvalues) is exponentially small. A similar phenomenon holds for topologically trivial spin glasses: those whose external fields are so strong that the landscape has only two critical points, which are the global extrema [BČNS22, XZ22, HS23c]. Topological trivialization is equivalent to the minimizing $\zeta$ being the constant function 00. In general, $\zeta$ can be much more complicated; see [AZ19, AZ22b] for examples. Aside from $1$-RSB and topologically trivial models, the annealed predictions from the Kac–Rice formula are incorrect, and do not suffice to understand the behavior near the ground state (though see [BJ24] for an application at positive temperature). In these cases, physicists have used the replica method to predict the correct quenched critical point counts; see the recent work [FL18, KK23] as well as [BM80, CS95]. These arguments seem difficult to make rigorous at present. Despite these challenges, it is still natural to believe that exponentially rare behavior does not occur at approximate ground states, because the sets $A_{\delta}$ must be rather small. Theorem 1.3 confirms this intuition for the absence of Hessian outliers, but uses a completely different proof technique.
The conclusion (1.9) is also related to thresholds obtained in [ABČ13, AB13] through the Kac–Rice formula. Namely in pure models $\xi(t)=t^{p}$, for marginal stability to hold at a critical point ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, one must have $H_{N}({\boldsymbol{\sigma}})/N\approx E_{\infty}(p)=2\sqrt{\frac{p-1}{p}}$. In mixed spherical spin glasses, a more involved Kac-Rice calculation shows that for marginally stable critical points to exist at energy $E$, one must have $E\in[E_{\infty}^{-},E_{\infty}^{+}]$ for
$E_{\infty}^{\pm}(\xi)\equiv\frac{2\xi^{\prime}\sqrt{\xi^{\prime\prime}}\pm\sqrt{4\xi^{\prime\prime}(\xi^{\prime})^{2}-\left(\xi^{\prime\prime}+\xi^{\prime}\right)\left(2\left(\xi^{\prime\prime}-\xi^{\prime}+(\xi^{\prime})^{2}\right)-\alpha^{2}\log\frac{\xi^{\prime\prime}}{\xi^{\prime}}\right)}}{\xi^{\prime}+\xi^{\prime\prime}}$
where $\alpha=\sqrt{\xi^{\prime\prime}+\xi^{\prime}-(\xi^{\prime})^{2}}$ and $\xi,\xi^{\prime},\xi^{\prime\prime}$ are evaluated at $1$. Of course, the exact ground state is a critical point of $H_{N}$. When $\xi$ has full RSB endpoint behavior, Proposition 1.2 implies it is marginally stable. Thus one obtains the following inequality, where the full RSB condition is necessary because ${\mathrm{GS}}(\xi)>E_{\infty}$ for pure models. By contrast ${\mathrm{GS}}(\xi)\geq E_{\infty}^{-}(\xi)$ holds for all $\xi$; see [AB13], or [HS23c, Section 7] for extensions.
Corollary 1.8.
If $\xi$ has full RSB endpoint behavior, then ${\mathrm{GS}}(\xi)\leq E_{\infty}^{+}(\xi)$.
1.3 On Marginal Stability
The type of qualitative connection between marginal stability and full RSB proved in Corollary 1.4 was long anticipated in the physics literature. For example, the Gardner transition [Gar85] from $1$-RSB to full RSB in Ising spin glasses is said to occur when the Hessian spectrum at ancestor states touches 00. We also quote from [MI04]:
[…]
See also [MW15, FU22]. It is separately understood, without clear mathematical justification, that “reasonable” optimization algorithms in high-dimension rapidly reach and then get stuck in the “manifold” of marginal states [CK94, Ken24]. As rigorous evidence for this, [HS24a] analyzed the optimal message-passing algorithms to optimize $H_{N}$ over ${\mathcal{S}}_{N}$ and showed the resulting outputs are marginally stable outside of the topologically trivial phase, even in the more general multi-species setting. Additionally, optimal stable algorithms to optimize mean-field spin glass Hamiltonians are now understood to be closely related to full RSB [Sub21, Mon21, AMS21, HS24b, AMS23b, HS23b, MZ24, JSS24]. Corollary 1.4 gives another rigorous link between full RSB and marginal stability, which however concerns low temperature statics rather than efficient algorithms.
1.4 Proof Ideas
The radial derivative formula (1.8) is proved using perturbative arguments in ${\mathbb{R}}^{N}$. The idea is to consider the maximum value of $H_{N}$ on a dilated spherical domain $t{\mathcal{S}}_{N}$. We take $|t-1|$ small (but independent of $N$), so the change in the maximum of $H_{N}$ is essentially given by the maximum or minimum radial derivative value at any approximate ground state (depending on the sign of $t-1$). On the other hand, it follows from (1.2) that the asymptotic ground state energy on a dilated sphere is given by ${\mathrm{GS}}(\xi_{t})$ for some slightly perturbed $\xi_{t}\approx\xi$. In this way, our proof of (1.8) first establishes an abstract formula for the radial derivative at any approximate ground state, expressed as a derivative of ${\mathrm{GS}}(\xi_{t})$. This relies on continuous differentiability properties of the spherical Parisi formula, which stem from the strict convexity of ${\mathcal{Q}}$. We then substantially simplify this formula using the stationarity conditions for the minimizing $(\zeta,L)$, yielding (1.8) and hence also (1.9).
The proofs of (1.11) and (1.12), which ensure the absence of Hessian outliers, are more delicate and given in Subsections 2.2 and 2.3 respectively. Here we perturb $H_{N}$ by considering augmented Hamiltonians of two or three replicas with constrained overlaps very close to $1$. By construction, the near-extrema for the augmented systems are obtained by starting from ${\boldsymbol{\sigma}}\in A_{\delta}$, and perturbing ${\boldsymbol{\sigma}}$ in opposite directions along a Hessian eigenvector. Thus in each case, precise estimates on the augmented ground state energy translate to bounds on the extreme eigenvalues at any approximate ground state. We find explicit interpolation parameters in the corresponding vector Parisi formulas, and perform a first-order Taylor expansion of the resulting bounds for the augmented ground state energy. This yields eigenvalue estimates which turn out to match the edge of the bulk spectrum from (1.9). Here $\xi$ is required to be even so the interpolation bounds for vector spin glasses are valid.
Finally Section 3 gives an alternate proof that Gibbs samples are marginally stable if $\beta\to\infty$ slowly with $N$, if $\xi$ is generic in addition to exhibiting full RSB endpoint behavior. Unlike the main proof described above, this alternate proof does not require any computations with the Parisi formula. Instead the full RSB condition is used to ensure that low-temperature Gibbs samples have positive probability to form certain ultrametric constellations with overlap close to $1$. Existence of such constellations then forces the Hessian to have many near-zero eigenvalues. This provides a more conceptual explanation for the link between marginal stability and full RSB.
1.5 Technical Preliminaries
Here we explicitly define the relevant derivative operations on the sphere, and state some useful concentration estimates. First, the rescaled radial derivative at ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$ is
$\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})=R({\boldsymbol{\sigma}},\nabla H_{N}({\boldsymbol{\sigma}})).$
Next for each ${\boldsymbol{\sigma}}\in S_{N}$, let $\{e_{1}({\boldsymbol{\sigma}}),\ldots,e_{N}({\boldsymbol{\sigma}})\}$ be an orthonormal basis of ${\mathbb{R}}^{N}$ with $e_{1}({\boldsymbol{\sigma}})={\boldsymbol{\sigma}}/\sqrt{N}$. Let ${\mathcal{T}}=\{2,\ldots,N\}$. Let $\nabla_{\mathcal{T}}H_{N}({\boldsymbol{\sigma}})\in{\mathbb{R}}^{\mathcal{T}}$ denote the projection of $\nabla H_{N}({\boldsymbol{\sigma}})\in{\mathbb{R}}^{N}$ to the space spanned by $\{e_{2}({\boldsymbol{\sigma}}),\ldots,e_{N}({\boldsymbol{\sigma}})\}$, and $\nabla^{2}_{{\mathcal{T}}\times{\mathcal{T}}}H_{N}({\boldsymbol{\sigma}})\in{\mathbb{R}}^{{\mathcal{T}}\times{\mathcal{T}}}$ analogously. The spherical gradient and Hessian are defined by:
$\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})=\nabla_{{\mathcal{T}}}H_{N}({\boldsymbol{\sigma}}),\quad\quad\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})=\nabla^{2}_{{\mathcal{T}}\times{\mathcal{T}}}H_{N}({\boldsymbol{\sigma}})-\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})I_{{\mathcal{T}}\times{\mathcal{T}}}.$
The next proposition provides smoothness estimates for $H_{N}$, ensuring for example that the radial derivative has typical order $O(1)$, while the spherical gradient has typical norm $O(\sqrt{N})$. The operator norm of a tensor $\boldsymbol{A}\in({\mathbb{R}}^{N})^{\otimes k}$ is
${\|\boldsymbol{A}\|}_{\rm op}=\max_{\|{\boldsymbol{\sigma}}^{1}\|_{2},\ldots,\|{\boldsymbol{\sigma}}^{k}\|_{2}\leq 1}|\langle\boldsymbol{A},{\boldsymbol{\sigma}}^{1}\otimes\cdots\otimes{\boldsymbol{\sigma}}^{k}\rangle|\,.$
We denote by ${\mathscr{H}}_{N}\simeq\bigoplus_{p\geq 1}{\mathbb{R}}^{N^{p}}$ the set of mixed $p$-spin Hamiltonians on ${\mathcal{S}}_{N}$, which for fixed $\xi$ can be identified with the coefficients ${\boldsymbol{G}}^{(p)}$.
Proposition 1.9.
For fixed $\xi$ there exist constants $C,c>0$, and a sequence $(K_{N})_{N\geq 1}$ of sets $K_{N}\subseteq{\mathscr{H}}_{N}$, with:
  1. 1.
    ${\mathbb{P}}[H_{N}\in K_{N}]\geq 1-e^{-cN}$.
  2. 2.
    If $H_{N}\in K_{N}$ and ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathcal{S}}_{N}$, then
[…]
  3. 3.
    If $H_{N}\in K_{N}$ then
    $\left|{\mathrm{GS}}_{N}-{\mathrm{GS}}(\xi)\right|\leq\varepsilon.$ (1.19)
Proof.
[HS24b, Proposition 1.1] shows that (1.17) and (1.18) hold with probability $1-e^{-cN}$. The same holds for (1.19) by the Borell-TIS inequality. ∎
Corollary 1.10.
Let $\delta$ be small depending on $(\xi,\varepsilon)$. Then with probability $1-e^{-cN}$, the following hold for all ${\boldsymbol{\sigma}}\in A_{\delta}$:
$|H_{N}({\boldsymbol{\sigma}})/N-{\mathrm{GS}}_{N}(\xi)|$ $\leq\varepsilon,$ (1.20)
$\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}})\|$ $\leq\varepsilon\sqrt{N},$ (1.21)
${\boldsymbol{\lambda}}_{1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}$ $\leq\varepsilon.$ (1.22)
Proof.
The first bound (1.20) follows by (1.19). This implies (1.21) and (1.22) because if one of them did not hold, then for $H_{N}\in K_{N}$, Taylor expanding $H_{N}$ near ${\boldsymbol{\sigma}}$ would show (1.20) is false for some $\varepsilon^{\prime}$. ∎
The next standard lemma shows that $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})$ determines the bulk spectral edge of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$ uniformly over ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$. In particular, (1.9) and (1.8) are equivalent (the case $N(1-\eta)\leq j\leq N-K$ follows by negating $H_{N}$, which preserves its law). We note that in Section 3, it is important to take $j$ constant in Lemma 1.11.
Lemma 1.11 ([Sub21, Lemma 3]).
For any $\varepsilon>0$ there are $K=K(\xi,\varepsilon)$ and $\eta=\eta(\xi,\varepsilon)>0$ such that
$\big{|}{\boldsymbol{\lambda}}_{j}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})\big{)}-2\xi^{\prime\prime}(1)-\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})\big{|}\leq\varepsilon$
holds for all $K\leq j\leq\eta N$ and ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$ simultaneously, with probability $1-e^{-cN}$.
Finally we record the standard fact that low temperature Gibbs samples are approximate ground states.
Lemma 1.12.
Assume $H_{N}\in K_{N}$. Then $A_{\beta^{-1/2}}$ (recall (1.7)) satisfies $\mu_{\beta}(A_{\beta^{-1/2}})\geq 1-e^{-cN}$ for $\beta\geq\beta_{*}(\xi)$ large.
Proof.
Let ${\boldsymbol{\sigma}}_{*}\in{\mathcal{S}}_{N}$ be the maximizer of $H_{N}$. The radius $\beta^{-2}\sqrt{N}$ neighborhood of ${\boldsymbol{\sigma}}_{*}$ has $\mu_{0}$ (uniform) measure $\beta^{-O(N)}$ and energy within $O(\beta^{-2}N)$ of the maximum. The contribution to $Z_{N,\beta}$ from this neighborhood is
$\exp\Big{(}\beta H_{N}({\boldsymbol{\sigma}}_{*})-N\cdot O(\log\beta)\Big{)}.$
When $\beta$ is large, this is exponentially larger than the contribution to $Z_{N,\beta}$ from the complement of $A_{\beta^{-1/2}}$. ∎
1.6 Characterization of the Minimizer in the Parisi Formula
Since the Parisi functional ${\mathcal{Q}}$ is strictly convex, the unique minimizer $(\zeta,L)$ is characterized by first-order stationarity conditions. These will be useful later, and we review them now. Given $(\zeta,L)\in{\mathcal{K}}$, define
$G(q)$ $=\xi^{\prime}(q)-\int_{0}^{q}\frac{{\rm d}s}{\hat{\zeta}(s)^{2}},$ $g(s)$ $=\int_{s}^{1}G(q)\leavevmode\nobreak\ {\rm d}q.$ (1.23)
Let $\nu$ be the finite Borel measure on $[0,1]$ defined by
$\nu([0,q])=\zeta(q)\qquad\forall q\in[0,1]$ (1.24)
and define the set
$T=\{q\in[0,1]:g(q)=0\}.$ (1.25)
Note that $1\in T$ holds trivially.
The characterization below is primarily from [CS17, Theorem 2]. The fact that $T$ consists of finitely many intervals is proved in [JT18, Corollary 1.6] at finite temperature, and the proof is essentially the same. One combines [JT17, Theorem 1.13] and the observation that by analyticity, $\big{(}\frac{1}{\sqrt{\xi^{\prime\prime}}}\big{)}^{\prime\prime}$ changes sign finitely many times on $[0,1]$.
Proposition 1.13.
There exists a unique $(L,\zeta)$ attaining the infimum (1.4), which is characterized by the following. Here ${\rm supp}(\zeta)\subseteq[0,1)$ denotes the set of points of increase of $\zeta$, i.e. the support of $\nu$.
$G(1)=0;\quad\quad\min_{q\in[0,1]}g(q)=0;\quad\quad T\subseteq{\rm supp}(\zeta).$
Furthermore, $T$ is a disjoint union of finitely many closed intervals (possibly including singletons).
Corollary 1.14.
If $1$ is not an isolated point in $T$, then $\xi$ exhibits full RSB endpoint behavior.
Proof.
By the last assertion of Proposition 1.13, $T$ must contain a non-trivial interval $(1-\varepsilon(\xi),1)$. Then $g^{\prime\prime}(1)=0$ by definition of $G$, which implies (1.10). ∎
2 Proofs of Main Results
Here we prove Proposition 1.2, Theorem 1.3, and Corollary 1.7. Subsection 2.1 computes the radial derivative at approximate ground states, thus proving (1.8). As discussed previously, this is done by considering the maximum value of $H_{N}$ on slight dilations of ${\mathcal{S}}_{N}$, and then simplifying the resulting abstract formula using the stationarity conditions from Subsection 1.6. The next two subsections then respectively prove (1.11) and (1.12) using multi-replica interpolation bounds. Finally Subsection 2.4 explains how to deduce Corollary 1.7 from Corollary 1.5.
2.1 Radial Derivative at Approximate Ground States
We fix $\xi$ and $\eta<0.01$ with $\xi(1+\eta)<\infty$, and let $\xi_{t}(q)=\xi(t^{2}q)$ for $t\in(1-\eta,1+\eta)$. To explain this definition, note that by (1.1) or (1.2), the function on ${\mathcal{S}}_{N}$ defined by ${\boldsymbol{\sigma}}\to H_{N}(t{\boldsymbol{\sigma}})$ is precisely a spherical spin glass with mixture $\xi_{t}$.
We would like to show $t\mapsto{\mathrm{GS}}(\xi_{t})$ is continuously differentiable using the envelope theorem (see e.g. [DM24, Chapter 2.4]). The first step is to check the minimizer varies within a compact space; note that apriori the space ${\mathcal{K}}$ is not compact. Thus, let
$(\zeta_{t},L_{t})=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi_{t})$
be the corresponding minimizers in the zero-temperature Parisi formula.
Lemma 2.1.
There exists $C=C(\xi)>0$ such that for all $t\in(1-\eta,1+\eta)$,
$L_{t}\leq C;\quad\quad\hat{\zeta}_{t}(1)\geq 1/C.$ (2.1)
Proof.
We treat $C=C(\xi)$ as a constant that may vary from line to line. In the topologically trivial case that $\zeta_{t}$ is identically 00, we easily find $L_{t}=\sqrt{\xi_{t}^{\prime}(1)}$ and so $\hat{\zeta}_{t}(1)=1/\sqrt{\xi_{t}^{\prime}(1)}$, so the conclusion is obvious. We assume this is not the case below. Noting that $\hat{\zeta}_{t}(\cdot)$ is a decreasing function of $t$, it follows that $\hat{\zeta}_{t}(1/2)\leq C$ is bounded independently of $t$ since
$2{\mathrm{GS}}(\xi_{t})$ $\geq 2{\mathcal{Q}}(\zeta_{t},L_{t};\xi_{t})\geq\int_{0}^{1}\xi_{t}^{\prime\prime}(q)\hat{\zeta}(q){\rm d}q\geq\int_{0}^{1/2}\xi_{t}^{\prime\prime}(q)\hat{\zeta}(q){\rm d}q$
$\geq\frac{\hat{\zeta}_{t}(1/2)}{2}\int_{0}^{1/2}\xi_{t}^{\prime\prime}(q)\leavevmode\nobreak\ {\rm d}q\geq\frac{\hat{\zeta}_{t}(1/2)}{2}\big{(}\xi^{\prime}(1/2)-\xi^{\prime}(0)\big{)}\geq\Omega(\hat{\zeta}_{t}(1/2)).$
Therefore $\hat{\zeta}_{t}(q)\geq L_{t}-C$ for all $q\in[1/4,1/2]$. Since $\xi_{t}^{\prime\prime}(1/4)\geq 1/C$ for all $t\in[1-\eta,1+\eta]$, we find that $L_{t}\leq C$ is also bounded independently of $t$.
We now turn to the second estimate. Let $A=\xi_{1+\eta}^{\prime\prime}(1)^{-1/2}\leq\xi_{t}^{\prime\prime}(1)^{-1/2}$, and suppose without loss of generality that $\hat{\zeta}_{t}(1)<A$. Let $q_{*}\in[0,1)$ satisfy $\hat{\zeta}_{t}(q_{*})=A$, and consider
$\zeta(q)=\begin{cases}\zeta(q),\quad q\leq q_{*};\\ \zeta(q_{*}),\quad q\geq q_{*}.\end{cases}$
Then by definition $\hat{\zeta}(q)\in[\hat{\zeta}_{t}(q),A]$ for all $q\in[0,1]$, so if $\zeta\neq\zeta_{t}$, we easily find that ${\mathcal{Q}}(\zeta,L_{t})<{\mathcal{Q}}(\zeta_{t},L_{t})$ which is a contradiction. (By similar reasoning, if $\hat{\zeta}(0)<A$, then increasing $L_{t}$ would decrease ${\mathcal{Q}}$, so $q_{*}$ actually exists.) Therefore $\zeta_{t}(q)=z$ must be constant on $q\in[q_{*},1)$.
Next, let $q^{*}\leq q_{*}$ be the largest point in the support of $\zeta_{t}$, with $q^{*}=0$ if $\zeta_{t}$ is identically 00. Recalling the notation from (1.23), we have $G(1)=G(q^{*})$ which means
$\xi_{t}^{\prime}(1)-\xi_{t}^{\prime}(q^{*})=\int_{q^{*}}^{1}\frac{{\rm d}q}{\hat{\zeta}_{t}(q)^{2}}=\frac{1}{z\hat{\zeta}_{t}(1)}-\frac{1}{z\hat{\zeta}_{t}(q^{*})}.$
Note that $\hat{\zeta}_{t}(q^{*})\geq\hat{\zeta}_{t}(q_{*})=A$. If $\hat{\zeta}_{t}(1)\geq A/2$ we are done; if not, we conclude that
$\xi_{t}^{\prime}(1)-\xi_{t}^{\prime}(q^{*})\geq\frac{1}{2z\hat{\zeta}_{t}(1)}\implies\hat{\zeta}_{t}(1)\geq\frac{1}{(1-q^{*})zC^{\prime}(\xi)}.$
However by our assumptions and the first part above, $(1-q^{*})z=\hat{\zeta}_{t}(1)-\hat{\zeta}_{t}(q^{*})\leq L_{t}\leq C$. Thus $\hat{\zeta}_{t}(1)$ is bounded away from 00 depending only on $\xi$ in all cases. ∎
Given $C>0$, let ${\mathcal{K}}_{C}\subseteq{\mathcal{K}}$ consist of those $(\zeta,L)$ satisfying (2.1). We endow ${\mathcal{K}}_{C}$ with the weak*-topology from the associated $\nu$ in (1.24), making it a compact metric space for each $C$.
Proposition 2.2.
For $s\in(1-\eta,1+\eta)$, the function $s\mapsto{\mathrm{GS}}(\xi_{s})$ is continuously differentiable with derivative
$\frac{{\rm d}}{{\rm d}s}{\mathrm{GS}}(\xi_{s})=\frac{{\rm d}}{{\rm d}t}{\mathcal{Q}}(\zeta_{s},L_{s};\xi_{t})\big{|}_{t=s}=\xi_{s}^{\prime}(0)L_{s}+\int_{0}^{1}\big{(}2\xi_{s}^{\prime\prime}(q)+q\xi_{s}^{\prime\prime\prime}(q)\big{)}\hat{\zeta}_{s}(q)\leavevmode\nobreak\ {\rm d}q.$ (2.2)
Proof.
Lemma 2.1 shows ${\mathrm{GS}}(\xi_{s})$ is defined as an infimum over a fixed compact set ${\mathcal{K}}_{C}\subseteq{\mathcal{K}}$; Proposition 1.13 guarantees this minimizer is unique. This easily implies $s\mapsto(\zeta_{s},L_{t})$ is continuous. The first equality now follows from the envelope theorem [MS02, Theorem 2 and Corollary 4], assuming the latter derivative of $Q$ exists. Indeed dominated convergence shows it is given by the explicit formula above. ∎
Lemma 2.3.
$r(\xi_{s})$ agrees with the formula (2.2).
Proof.
We use the stationarity conditions reviewed in Subsection 1.6, and set $s=1$ for convenience. Since $\hat{\zeta}^{\prime}(q)=-\zeta(q)$, integrating by parts gives
$\int_{0}^{1}\big{(}\xi^{\prime\prime}(q)+q\xi^{\prime\prime\prime}(q)\big{)}\hat{\zeta}(q){\rm d}q=\underbrace{[q\xi^{\prime\prime}(q)\hat{\zeta}(q)]\big{|}_{0}^{1}}_{\xi^{\prime\prime}(1)\hat{\zeta}(1)}+\int_{0}^{1}q\xi^{\prime\prime}(q)\zeta(q){\rm d}q.$
Recalling the definition (1.5) of $r(\xi)$, it remains to show that
$\xi^{\prime}(0)L+\int_{0}^{1}\xi^{\prime\prime}(q)\hat{\zeta}(q)+q\xi^{\prime\prime}(q)\zeta(q){\rm d}q\stackrel{{\scriptstyle?}}{{=}}\hat{\zeta}(1)^{-1}.$ (2.3)
Note that by the last assertion in Proposition 1.13, there exists a unique finite sequence $10\leq q_{0}<q_{1}<\dots<q_{D}=$ 1 such that for each $0\leq d\leq D$, either one of the closed intervals $[q_{d-1},q_{d}]$ or $[q_{d},q_{d+1}]$ is a connected component of $T$, or $q_{d}$ is an isolated point. We verify (2.3) by evaluating the integral separately over each subinterval $[q_{d},q_{d+1}]$.
We first handle the contribution from non-trivial intervals $[q_{d},q_{d+1}]\subseteq T$. In this case, (1.23) gives $g^{\prime\prime}(q)=0$ for all $q\in[q_{d},q_{d+1}]$, implying $\hat{\zeta}(q)=\xi^{\prime\prime}(q)^{-1/2}$. Then $\zeta(q)=\frac{\xi^{\prime\prime\prime}(q)}{2\xi^{\prime\prime}(q)^{3/2}}$, so the contribution to the integral in (2.3) is:
$\int_{q_{d}}^{q_{d+1}}\xi^{\prime\prime}(q)^{1/2}+\frac{q\xi^{\prime\prime\prime}(q)}{2\xi^{\prime\prime}(q)^{1/2}}{\rm d}q=[q\xi^{\prime\prime}(q)^{1/2}]\big{|}_{q_{d}}^{q_{d+1}}=\frac{q_{d+1}}{\hat{\zeta}(q_{d+1})}-\frac{q_{d}}{\hat{\zeta}(q_{d})}.$
Next suppose $[q_{d},q_{d+1}]\cap T=\{q_{d},q_{d+1}\}$. Then $\zeta$ is constant on $(q_{d},q_{d+1})$, and so on this interval $\hat{\zeta}(q)=x-yq$ for some real $(x,y)=(x_{d}(\xi),y_{d}(\xi))$ with $y\geq 0$, where $\zeta(q)=y$. Then the contribution is
$\int_{q_{d}}^{q_{d+1}}\xi^{\prime\prime}(q)(x-yq)+q\xi^{\prime\prime}(q)y{\rm d}q=x\big{(}\xi^{\prime}(q_{d+1})-\xi^{\prime}(q_{d})\big{)}.$
Since $g$ is minimized at both $q_{d}$ and $q_{d+1}$, we have $G(q_{d})=G(q_{d+1})=0$. Thus we similarly find:
$x\big{(}\xi^{\prime}(q_{d+1})-\xi^{\prime}(q_{d})\big{)}$ $=x\int_{q_{d}}^{q_{d+1}}\frac{{\rm d}q}{\hat{\zeta}(q)^{2}}=\int_{q_{d}}^{q_{d+1}}\frac{x}{(x-yq)^{2}}{\rm d}q$
$=\left[\frac{q}{x-yq}\right]\Big{|}_{q_{d}}^{q_{d+1}}=\frac{q_{d+1}}{\hat{\zeta}(q_{d+1})}-\frac{q_{d}}{\hat{\zeta}(q_{d})}.$
Telescoping,
$\int_{q_{0}}^{1}\xi^{\prime\prime}(q)\hat{\zeta}(q)+q\xi^{\prime\prime}(q)\zeta(q){\rm d}q=\frac{1}{\hat{\zeta}(1)}-\frac{q_{0}}{\hat{\zeta}(q_{0})}.$
If $q_{0}=0$, this completes the proof (this condition is equivalent to $\xi^{\prime}(0)=0$). If not, since $\zeta(q)=0$ for $q<q_{0}$, the function $\hat{\zeta}$ is constant on $[0,q_{0}]$ and so
$\int_{0}^{q_{0}}\xi^{\prime\prime}(q)\hat{\zeta}(q)\leavevmode\nobreak\ {\rm d}q=L\big{(}\xi^{\prime}(q_{0})-\xi^{\prime}(0)\big{)}.$
Furthermore $G(q_{0})=0$, since either $q_{0}\in(0,1)$ must be an interior minimizer of $g$ (which implies $G(q_{0})=-g^{\prime}(q_{0})=0$), or $q_{0}=1$ so again $G(q_{0})=0$ by Proposition 1.13. Thus
$\xi^{\prime}(q_{0})=\int_{0}^{q_{0}}\frac{{\rm d}q}{\hat{\zeta}(q)^{2}}=\frac{q_{0}}{\hat{\zeta}(0)^{2}}=\frac{q_{0}}{L^{2}}.$
Thus the first term in (2.3) is $\xi^{\prime}(0)L=\frac{q_{0}}{L}=\frac{q_{0}}{\hat{\zeta}(q_{0})}$ when $q_{0}>0$. Matching terms finishes the proof. ∎
Next we apply Proposition 2.2 and standard concentration estimates to determine the radial derivative at any near-ground state, thus proving (1.8). The point is that, as explained at the start of this subsection, if $\xi$ is the mixture function for $H_{N}({\boldsymbol{\sigma}})$, then $\xi_{t}$ is the mixture function for $H_{N}(t{\boldsymbol{\sigma}})$. On the other hand, elementary calculus in ${\mathbb{R}}^{N}$ shows the radial derivative at ground states determines the change in the maximum value of $H_{N}$ from a small dilation of ${\mathcal{S}}_{N}$.
Proof of Proposition 1.2.
Due to Lemma 1.11, it suffices to prove (1.8). In light of (1.21) in Corollary 1.10, the nontrivial portion of (1.8) is to compute the radial derivative. We restrict to the event $K_{N}$ of Proposition 1.9. Let $H_{N}$ be a random Hamiltonian with mixture $\xi$ and let
${\mathrm{GS}}(H_{N};t)=\max_{{\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}}H_{N}(t{\boldsymbol{\sigma}})/N.$
Setting $\eta=\sqrt{\delta}$, we consider the event that
$|{\mathrm{GS}}(H_{N};t)-{\mathrm{GS}}(\xi_{t})|\leq\delta,\quad\quad\forall\,t\in\{1-\eta,1,1+\eta\}.$ (2.4)
It follows from the discussion above, e.g. (1.2), that this event has probability at least $1-e^{-cN}$. We will show it implies the desired conclusion.
Assume for sake of contradiction that $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})\geq r(\xi)+\varepsilon$ for some $\delta$-approximate ground state ${\boldsymbol{\sigma}}$; the opposite case is similar using $1-\eta$ instead of $1+\eta$. Taylor expanding and bounding the error terms with Proposition 1.9,
$\frac{H_{N}((1+\eta){\boldsymbol{\sigma}})-H_{N}({\boldsymbol{\sigma}})}{N}\geq\eta\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})-C(\xi)\eta^{2}.$
Since ${\boldsymbol{\sigma}}$ is an $\eta^{2}$-approximate ground state for $H_{N}$, we find from (2.4) and the assumption on $\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})$ that
${\mathrm{GS}}(\xi_{1+\eta})$ $\geq H_{N}((1+\eta){\boldsymbol{\sigma}})/N-\delta$
$\geq H_{N}({\boldsymbol{\sigma}})/N+\eta\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})-(C(\xi)+1)\eta^{2}$
$\geq{\mathrm{GS}}(\xi_{1})+\eta\partial_{{\mathrm{rad}}}H_{N}({\boldsymbol{\sigma}})-(C(\xi)+2)\eta^{2}$
$\geq{\mathrm{GS}}(\xi_{1})+\eta r(\xi)+\eta\varepsilon-(C(\xi)+2)\eta^{2}.$
On the other hand Proposition 2.2 ensures the continuous differentiability of ${\mathrm{GS}}(\xi_{t})$, so for $\eta\leq\eta_{*}(\xi,\varepsilon)$ small,
${\mathrm{GS}}(\xi_{1+\eta})\leq{\mathrm{GS}}(\xi_{1})+\eta r(\xi)+\frac{\eta\varepsilon}{2}.$
The previous two displays are contradictory for $\eta$ small depending on $(\xi,\varepsilon)$, which completes the proof. ∎
2.2 No Upward Outliers in Even Models
When $\xi(q)=\sum_{p\in 2{\mathbb{N}}}\gamma_{p}^{2}q^{p}$ is even, we show that the Hessian at near ground states has no outliers. We first show (1.11), the absence of eigenvalues above the bulk, by estimating the ground state energy of a two-replica spin glass:
${\mathrm{GS}}_{2,\varepsilon}(H_{N})=\sup_{\begin{subarray}{c}{\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in{\mathcal{S}}_{N}\\ R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})=1-\varepsilon\end{subarray}}\frac{H_{N}({\boldsymbol{\sigma}})+H_{N}({\boldsymbol{\sigma}}^{\prime})}{N}.$
When $\xi$ is even, the limit ${\mathrm{GS}}_{2,\varepsilon}(\xi)=\operatorname*{p-lim}_{N\to\infty}{\mathrm{GS}}_{2,\varepsilon}(H_{N})$ exists and is given by a two-dimensional constrained generalization of the Parisi formula. Although we refrain from giving the notationally heavy general definitions, vector models of this type have played an important role in spin glass theory, being used in Talagrand’s original proof of the Parisi formula [Tal06b, Tal06a] and subsequent works [PT07, CHHS15, Che17, CHL18, AC18, CGPR19].
The multi-replica Parisi formula is given by a similar variational problem as in Proposition 1.1. We choose suitable interpolation parameters to upper bound ${\mathrm{GS}}_{2,\varepsilon}(\xi)$ for small $\varepsilon$, and thus deduce non-existence of outlier eigenvalues at approximate ground states. The general formula replaces $L$ by a $2\times 2$ positive semi-definite matrix $\boldsymbol{L}$, and $\zeta$ by a cumulative distribution function on a monotone path of $2\times 2$ matrices in the positive semi-definite order. The path of matrices is encapsulated by an function $\Phi:[0,2]\to{\mathbb{R}}^{2\times 2}$ with $\Phi(0)=0,\Phi(2)=Q$ and ${\sf Tr}(\Phi(t))=t$ for all $t$, with $\Phi(t)-\Phi(s)$ positive semi-definite for all $t\geq s$. Meanwhile $\alpha:[0,1]\to{\mathbb{R}}_{+}$ is the associated cumulative distribution function, which can be viewed as a positive measure on the range of $\Phi$. For us, the relevant specialization is as follows. Define the $2\times 2$ matrices:
$J_{+}=\begin{pmatrix}1&1\\ 1&1\end{pmatrix};\quad\quad J_{-}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix};\quad\quad Q=\begin{pmatrix}1&1-\varepsilon\\ 1-\varepsilon&1\end{pmatrix}=\Big{(}1-\frac{\varepsilon}{2}\Big{)}J_{+}+\frac{\varepsilon J_{-}}{2}.$
We take as given $(\zeta,L)=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi)$. With $\ell=\xi^{\prime\prime}(1)^{-1/2}$, their two-replica generalizations are:
$\boldsymbol{L}=\frac{LJ_{+}+\ell J_{-}}{2};\quad\quad\alpha(t)=\zeta(t/2)/2;\quad\quad\Phi(t)=\begin{cases}\frac{tJ_{+}}{2},\quad t\in[0,2-\varepsilon];\\ \frac{(2-\varepsilon)J_{+}}{2}+\frac{t-(2-\varepsilon)J_{-}}{2},\quad t\in[2-\varepsilon,2].\end{cases}$
Applying [AZ22a, Theorem 6] with $(L,\alpha,\Phi)$ as above gives the following upper bound on ${\mathrm{GS}}_{2,\varepsilon}$ (see also [Ko20]). Then Corollary 2.5 gives a first order expansion of this bound for small $\varepsilon$, which turns out to be sharp.
Proposition 2.4.
Let $\xi$ be even, with minimizer $(\zeta,L)=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi)$. Then with $\xi,\xi^{\prime},\xi^{\prime\prime}$ acting entry-wise on $2\times 2$ matrices, $\langle A,B\rangle={\sf Tr}(A^{\top}B)$, and $\odot$ denoting entry-wise product:
$2\,{\mathrm{GS}}_{2,\varepsilon}(\xi)$ $\leq\langle\xi^{\prime}(Q),\boldsymbol{L}\rangle-\int_{0}^{2}\Big{\langle}\xi^{\prime\prime}(\Phi(t))\odot\Phi^{\prime}(t),\int_{0}^{t}\alpha(s)\Phi^{\prime}(s)\,{\rm d}s\rangle\,{\rm d}t\Big{\rangle}$
$\quad\quad+\int_{0}^{2}\Big{\langle}\big{(}\boldsymbol{L}-\int_{0}^{t}\alpha(s)\Phi^{\prime}(s){\rm d}s\big{)}^{-1},\Phi^{\prime}(t)\Big{\rangle}\,{\rm d}t.$
Corollary 2.5.
For any $\iota>0$ there exists $\varepsilon_{*}=\varepsilon_{*}(\xi,\iota)>0$ such that for all $\varepsilon\in(0,\varepsilon_{*})$:
$\frac{{\mathrm{GS}}_{2,\varepsilon}(\xi)-2{\mathrm{GS}}(\xi)}{\varepsilon}-\iota\leq\sqrt{\xi^{\prime\prime}(1)}-\frac{r(\xi)}{2}=-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}.$
Proof.
We evaluate the upper bound for $2{\mathrm{GS}}_{2,\varepsilon}(\xi)$ in Proposition 2.4 to first order in $\varepsilon$. The first term is
$L\big{(}\xi^{\prime}(1)+(1-\varepsilon)\xi^{\prime}(1)\big{)}=2L\xi^{\prime}(1)-L\xi^{\prime\prime}(1)\varepsilon+\ell\xi^{\prime\prime}(1)\varepsilon+O(\varepsilon^{2}).$
Via the linear changes of variable $(u,v)=(s/2,t/2)$, the integral contribution from $[0,2-\varepsilon]$ in the second term is:
$\int\limits_{0}^{2-\varepsilon}\left\langle\frac{\xi^{\prime\prime}(t/2)J_{+}}{2},J_{+}\int_{0}^{t}\zeta(s/2)/4\leavevmode\nobreak\ {\rm d}s\right\rangle{\rm d}t$ $=2\int\limits_{0}^{1-\frac{\varepsilon}{2}}\xi^{\prime\prime}(r)\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right){\rm d}v$
$=2\int_{0}^{1}\xi^{\prime\prime}(r)\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right){\rm d}v-\left(\xi^{\prime\prime}(1)\int_{0}^{1}\zeta(u)\,{\rm d}u\right)\varepsilon+O(\varepsilon^{2}).$
The integral contribution from $[2-\varepsilon,2]$ in the second term is $O(\varepsilon^{2})$, since the integrand takes the form
$\Big{\langle}O(\varepsilon)J_{+}+\Theta(1)J_{-},\leavevmode\nobreak\ \Theta(1)J_{+}+O(\varepsilon)J_{-}\Big{\rangle}=O(\varepsilon).$
For the third term, we have $\Phi^{\prime}(t)=J_{+}/2$ on $t\in[0,2-\varepsilon]$. Note that
$(AJ_{+}+BJ_{-})^{-1}=(A^{-1}J_{+}+B^{-1}J_{-})/4,\quad\forall A,B\neq 0.$ (2.5)
Hence we only need to track the $J_{+}$ component in the matrix $\boldsymbol{L}-\int_{0}^{t}\alpha(s)\Phi^{\prime}(s){\rm d}s$ to determine the integral contribution on $[0,2-\varepsilon]$. This yields:
$\int_{0}^{2-\varepsilon}\left(\frac{L}{2}-\frac{1}{2}\int_{0}^{t/2}\zeta(u){\rm d}u\right)^{-1}\frac{1}{2}\leavevmode\nobreak\ {\rm d}t$ $=\int_{0}^{2-\varepsilon}\frac{{\rm d}t}{L-\int_{0}^{t/2}\zeta(u){\rm d}u}$
$=2\int_{0}^{1}\frac{{\rm d}q}{L-\int_{0}^{q}\zeta(u){\rm d}u}-\frac{\varepsilon}{\hat{\zeta}(1)}+O(\varepsilon^{2}).$
Finally the contribution on $[2-\varepsilon,2]$ is
$\int_{2-\varepsilon}^{2}\Big{(}\ell-\int_{2-\varepsilon}^{t}\zeta(s/2)\,{\rm d}s\Big{)}^{-1}{\rm d}t=\int_{1-\frac{\varepsilon}{2}}^{1}\Big{(}\frac{\ell}{2}-\int_{1-\frac{\varepsilon}{2}}^{v}\zeta(u)\,{\rm d}u\Big{)}^{-1}{\rm d}v=\frac{\varepsilon}{\ell}+O(\varepsilon^{2}).$
Recall that $\ell=\xi^{\prime\prime}(1)^{-1/2}$. Combining terms (and recalling that the second term is negated) yields
$2{\mathrm{GS}}_{2,\varepsilon}(\xi)\leq 4{\mathrm{GS}}(\xi)-\hat{\zeta}(1)\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+O(\varepsilon^{2}).\qed$
Using Corollary 2.5, we now deduce the “no upward outliers” property (1.11) for even models in Proposition 1.2. The idea is that any outlier eigenvalue yields a counterexample to the estimate have we just established.
Proof of Proposition 1.2, Eq. (1.11).
We restrict to the event of Proposition 1.9. Take $\varepsilon$ small depending on $(\xi,\iota)$, and $\delta\leq\varepsilon^{2}$. We consider the event that
$|{\mathrm{GS}}(H_{N})-{\mathrm{GS}}(\xi)|\leq\delta\quad\quad\text{and}\quad\quad{\mathrm{GS}}_{2,\varepsilon}(H_{N})\leq 2{\mathrm{GS}}(\xi)-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+\iota^{2}\varepsilon.$ (2.6)
This event clearly has probability at least $1-e^{-cN}$, and we will show it implies the desired conclusion. Let ${\boldsymbol{\sigma}}$ be a $\delta$-approximate ground state, and suppose for sake of contradiction that
${\boldsymbol{\lambda}}_{1}(\nabla_{\mathrm{sp}}^{2}(H_{N}({\boldsymbol{\sigma}})))\geq-\hat{\zeta}(1)\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}+\iota.$
(The lower bound on ${\boldsymbol{\lambda}}_{1}$ is trivial since ${\boldsymbol{\lambda}}_{1}(\cdot)\geq{\boldsymbol{\lambda}}_{\lfloor\delta N\rfloor}(\cdot)$.) Let $\theta(\varepsilon)=\arcsin(\sqrt{\varepsilon/2})=\sqrt{\varepsilon/2}+O(\varepsilon)$. With ${\boldsymbol{v}}/\sqrt{N}$ the maximum unit eigenvector of $\nabla_{\mathrm{sp}}^{2}(H_{N}({\boldsymbol{\sigma}}))$, set
${\boldsymbol{\sigma}}^{\pm}=\cos(\theta(\varepsilon)){\boldsymbol{\sigma}}\pm\sin(\theta(\varepsilon)){\boldsymbol{v}}.$ (2.7)
Then $R({\boldsymbol{\sigma}}^{+},{\boldsymbol{\sigma}}^{-})=\cos(2\theta(\varepsilon))=1-\varepsilon$. Since $\varepsilon\ll\iota$, Taylor expanding (via Proposition 1.9) gives
$\frac{H_{N}({\boldsymbol{\sigma}}^{+})+H_{N}({\boldsymbol{\sigma}}^{-})-2H_{N}({\boldsymbol{\sigma}})}{N}$ $=\frac{\varepsilon}{2}\cdot{\boldsymbol{\lambda}}_{1}(\nabla_{\mathrm{sp}}^{2}(H_{N}({\boldsymbol{\sigma}})))+O(\varepsilon^{2})$
$\geq-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+\Omega(\iota\varepsilon).$
However, (2.6) implies the left-hand expression is at most
$-\frac{\hat{\zeta}(1)}{2}\left(\sqrt{\xi^{\prime\prime}(1)}-\hat{\zeta}(1)^{-1}\right)^{2}\varepsilon+\iota^{2}\varepsilon.$
This is a contradiction, completing the proof (with $\iota$ here functioning as $\varepsilon$ in the original statement). ∎
2.3 No Downward Outliers in Even Models
The opposite estimate (1.12) ruling out downward outliers is conceptually similar but requires $3$ replicas. We consider the ground state energy of a three-replica Hamiltonian:
${\mathrm{GS}}_{3,\varepsilon}=\sup_{\begin{subarray}{c}{\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2},{\boldsymbol{\sigma}}^{3}\in{\mathcal{S}}_{N}\\ R({\boldsymbol{\sigma}}^{i},{\boldsymbol{\sigma}}^{j})=Q_{3}[i,j]\forall i,j\end{subarray}}\frac{H_{N}^{(3)}({\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2},{\boldsymbol{\sigma}}^{3})}{N}\equiv\sup_{\begin{subarray}{c}{\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2},{\boldsymbol{\sigma}}^{3}\in{\mathcal{S}}_{N}\\ R({\boldsymbol{\sigma}}^{i},{\boldsymbol{\sigma}}^{j})=Q_{3}[i,j]\forall i,j\end{subarray}}\frac{3H_{N}({\boldsymbol{\sigma}}^{1})-H_{N}({\boldsymbol{\sigma}}^{2})-H_{N}({\boldsymbol{\sigma}}^{3})}{N}.$ (2.8)
Here the $3\times 3$ overlap matrix $Q_{3}$ is given by:
$Q_{3}=\begin{pmatrix}1&1-\varepsilon&1-\varepsilon\\ 1-\varepsilon&1&1-4\varepsilon+2\varepsilon^{2}\\ 1-\varepsilon&1-4\varepsilon+2\varepsilon^{2}&1\end{pmatrix}\,.$
$Q_{3}$ is constructed to correspond to triples $({\boldsymbol{\sigma}}^{1},{\boldsymbol{\sigma}}^{2},{\boldsymbol{\sigma}}^{3})\in{\mathcal{S}}_{N}$ for which ${\boldsymbol{\sigma}}^{1}$ is on the midpoint of the geodesic between ${\boldsymbol{\sigma}}^{2}$ and ${\boldsymbol{\sigma}}^{3}$; note that if $\cos(\theta)=1-\varepsilon$, then $\cos(2\theta)=1-4\varepsilon+2\varepsilon^{2}$.
We again use interpolation to upper bound the limiting value ${\mathrm{GS}}_{3,\varepsilon}$, which now controls the minimum Hessian eigenvalue at an approximate ground state. The covariance structure of the Hamiltonian $H_{N}^{(3)}$ is less symmetric due to the coefficients $3$ and $-1$. Instead of extending $\xi$ by entry-wise application as before, we define ${\boldsymbol{\xi}}:{\mathbb{R}}^{3\times 3}\to{\mathbb{R}}^{3\times 3}$ by:
${\boldsymbol{\xi}}\begin{pmatrix}a_{1,1}&a_{1,2}&a_{1,3}\\ a_{2,1}&a_{2,2}&a_{2,3}\\ a_{3,1}&a_{3,2})&a_{3,3}\end{pmatrix}=\begin{pmatrix}9\xi(a_{1,1})&-3\xi(a_{1,2})&-3\xi(a_{1,3})\\ -3\xi(a_{2,1})&\xi(a_{2,2})&\xi(a_{2,3})\\ -3\xi(a_{3,1})&\xi(a_{3,2})&\xi(a_{3,3})\end{pmatrix}.$
This ${\boldsymbol{\xi}}$ is defined in general by the covariance structure induced by $H_{N}^{(3)}$ (recall (1.2)). Next, define the $3\times 3$ matrices:
$J_{3}=\begin{pmatrix}1&1-\varepsilon&1-\varepsilon\\ 1-\varepsilon&(1-\varepsilon)^{2}&(1-\varepsilon)^{2}\\ 1-\varepsilon&(1-\varepsilon)^{2}&(1-\varepsilon)^{2}\end{pmatrix};\quad J_{-}=\begin{pmatrix}0&0&0\\ 0&1&-1\\ 0&-1&1\end{pmatrix};\quad J_{*}=\begin{pmatrix}(2-2\varepsilon)^{2}&-(2-2\varepsilon)&-(2-2\varepsilon)\\ -(2-2\varepsilon)&1&1\\ -(2-2\varepsilon)&1&1\end{pmatrix}.$
Note that $(J_{3},J_{-},J_{*})=(v_{3}^{\otimes 2},v_{-}^{\otimes 2},v_{*}^{\otimes 2})$ for orthogonal vectors $(v_{3},v_{-},v_{*})$. This implies that, similarly to (2.5):
${\sf Tr}\big{(}(AJ_{3}+BJ_{-}+CJ_{*})^{-1}\cdot(DJ_{3}+EJ_{-}+FJ_{*})\big{)}={\sf Tr}\left(\frac{DJ_{3}}{A{\sf Tr}(J_{3})}+\frac{EJ_{-}}{B{\sf Tr}(J_{-}}+\frac{FJ_{*}}{C{\sf Tr}(J_{*})}\right)=\frac{D}{A}+\frac{E}{B}+\frac{F}{C}.$ (2.9)
We again fix $(\zeta,L)=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi)$ and take $\ell=\xi^{\prime\prime}(1)^{-1/2}$. This time, we set
$\boldsymbol{L}_{3}=LJ_{3}+\frac{\ell J_{-}}{2}+\varepsilon^{2}J_{*}.$
We note that the sole use of the matrix $J_{*}$ is to slightly increase $\boldsymbol{L}_{3}$ so the $3\times 3$ matrix inverses appearing below are well-defined; it will play no further role below.
Let $t_{*}=3-4\varepsilon+4\varepsilon^{2}={\sf Tr}(J_{3})$. We take $\Phi_{3}:[0,3]\to{\mathbb{R}}^{3\times 3}$ to be piece-wise linear with $\Phi_{3}(0)=0$, and
$\Phi_{3}^{\prime}(t)=\begin{cases}\frac{J_{3}}{t_{*}},\quad t\in[0,t_{*});\\ \frac{J_{-}}{2},\quad t\in(t_{*},3].\end{cases}$
It is easy to check that ${\sf Tr}(\Phi_{3}(t))=t$ for all $t\in[0,3]$, and $\Phi_{3}(3)=Q_{3}$. Further, $\Phi$ is again increasing in the positive semi-definite order. This time we set $\alpha_{3}(t)=\zeta\left(\frac{t}{3-t_{*}}\right)$ for $t\in[0,t_{*}]$, and extend $\alpha(t)$ to be constant on $t_{*},3]$. The following three-replica interpolation bound again follows directly from [AZ22b, Theorem 6].
Proposition 2.6.
Let $\xi$ be even, with minimizer $(\zeta,L)=\operatorname*{\rm arg\,min}_{(\zeta,L)\in{\mathcal{K}}}{\mathcal{Q}}(\zeta,L;\xi)$. Then
$2\,{\mathrm{GS}}_{3,\varepsilon}(\xi)$ $\leq\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),\boldsymbol{L}_{3}\rangle-\int_{0}^{3}\Big{\langle}{\boldsymbol{\xi}}^{\prime\prime}(\Phi_{3}(t))\odot\Phi_{3}^{\prime}(t),\int_{0}^{t}\alpha_{3}(s)\Phi_{3}^{\prime}(s)\,{\rm d}s\rangle\,{\rm d}t\Big{\rangle}$
$\quad\quad+\int_{0}^{3}\Big{\langle}\big{(}\boldsymbol{L}_{3}-\int_{0}^{t}\alpha_{3}(s)\Phi_{3}^{\prime}(s){\rm d}s\big{)}^{-1},\Phi_{3}^{\prime}(t)\Big{\rangle}\,{\rm d}t.$
The corresponding first-order expansion for this interpolation bound is as follows.
Corollary 2.7.
For any $\iota>0$ there exists $\varepsilon_{*}=\varepsilon_{*}(\xi,\iota)>0$ such that for all $\varepsilon\in(0,\varepsilon_{*})$:
$\frac{{\mathrm{GS}}_{3,\varepsilon}(\xi)-{\mathrm{GS}}(\xi)}{\varepsilon}-\iota\leq 2\hat{\zeta}(1)\left(\sqrt{\xi^{\prime\prime}(1)}+\hat{\zeta}(1)^{-1}\right)^{2}.$
Proof.
We evaluate the interpolation bound in Proposition 2.6 to first order in $\varepsilon$, writing $\approx$ for equalities up to $O(\varepsilon^{2})$ error. The first term is
$\left\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),\boldsymbol{L}_{3}\right\rangle\approx L\left\langle{\boldsymbol{\xi}}^{\prime}(Q_{3}),J_{3}\right\rangle$
$\quad\approx L\big{(}9\xi^{\prime}(1)-12(1-\varepsilon)\xi^{\prime}(1-\varepsilon)+2(1-2\varepsilon)\xi^{\prime}(1)+2(1-2\varepsilon)\xi^{\prime}(1-4\varepsilon)\big{)}+\frac{\ell}{2}\big{(}2\xi^{\prime}(1)-2\xi^{\prime}(1-4\varepsilon)\big{)}$
$\quad=L\xi^{\prime}(1)+4L\xi^{\prime}(1)\varepsilon+4L\xi^{\prime\prime}(1)\varepsilon+4\ell\xi^{\prime\prime}(1)\varepsilon+O(\varepsilon^{2}).$
The integral contribution from $[0,t_{*}]$ in the second term is, up to error $O(\varepsilon^{2})$:
$\frac{1}{t_{*}^{2}}\int\limits_{0}^{t_{*}}\left(\int_{0}^{t}\alpha_{3}(s)\,{\rm d}s\right)\left[9\xi^{\prime\prime}\left(\frac{t}{3-4\varepsilon}\right)-12\xi^{\prime\prime}\left(\frac{t}{3-\varepsilon}\right)(1-2\varepsilon)+4\xi^{\prime\prime}\left(\frac{t}{3+2\varepsilon}\right)(1-4\varepsilon)\right]{\rm d}t.$
Substituting $(u,v)=(s/t_{*},t/t_{*})$ and linearly approximating $\xi^{\prime\prime}$ near $t/t_{*}$, this approximately equals:
$\int_{0}^{1}\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right)\Big{[}9\xi^{\prime\prime}(v)-12(1-2\varepsilon)\xi^{\prime\prime}(v-\varepsilon v)+4(1-4\varepsilon)\xi^{\prime\prime}(v-2\varepsilon v)\Big{]}{\rm d}v$
$\approx\int_{0}^{1}\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right)\Big{[}\xi^{\prime\prime}(v)+\Big{(}8\xi^{\prime\prime}(v)+4v\xi^{\prime\prime\prime}(v)\Big{)}\varepsilon\Big{]}{\rm d}v.$
Recalling the main computation in Proposition 2.2, we find
$\int_{0}^{1}\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right)\big{(}2\xi^{\prime\prime}(v)+v\xi^{\prime\prime\prime}(v)\big{)}\,{\rm d}v$ $=L\underbrace{\int_{0}^{1}\big{(}2\xi^{\prime\prime}(v)+v\xi^{\prime\prime\prime}(v)\big{)}\,{\rm d}v}_{[\xi^{\prime}(v)+v\xi^{\prime\prime}(v)]|_{0}^{1}}-\int_{0}^{1}\hat{\zeta}(v)\big{(}2\xi^{\prime\prime}(v)+v\xi^{\prime\prime\prime}(v)\big{)}\,{\rm d}v$
$=L\xi^{\prime}(1)+L\xi^{\prime\prime}(1)-\Big{[}L\xi^{\prime}(0)+\int_{0}^{1}\hat{\zeta}(v)\big{(}2\xi^{\prime\prime}(v)+v\xi^{\prime\prime\prime}(v)\big{)}\,{\rm d}v\Big{]}$
[…]
The integral contribution from $[t_{*},3]$ in the second term is again $O(\varepsilon^{2})$, for similar reasons as before. Hence overall, the second term is
$\int_{0}^{1}\left(\int_{0}^{v}\zeta(u)\,{\rm d}u\right)\xi^{\prime\prime}(v)\,{\rm d}v+4\varepsilon\Big{[}L\xi^{\prime}(1)+L\xi^{\prime\prime}(1)-\hat{\zeta}(1)\xi^{\prime\prime}(1)-\hat{\zeta}(1)^{-1}\Big{]}+O(\varepsilon^{2}).$
The contribution to the third term from $[0,t_{*}]$ is made convenient by (2.9) and our choice of $\alpha_{3}$:
$\frac{1}{t_{*}}\int_{0}^{t_{*}}\left(L-\left(\int_{0}^{t/t_{*}}\zeta(u)\,{\rm d}u\right)\right)^{-1}{\rm d}t\approx\int_{0}^{1}\hat{\zeta}(v)^{-1}\,{\rm d}v..$
Finally on $[t_{*},3]$, the integrand is constant up to $O(\varepsilon)$ error, and $3-t_{*}\approx 4\varepsilon$. Hence the contribution is:
$4\varepsilon\left\langle\left(\boldsymbol{L}_{3}-\int_{0}^{3}\alpha_{3}(s)\Phi_{3}^{\prime}(s)\,{\rm d}s\right),J_{-}/2\right\rangle\approx\frac{4\varepsilon}{\ell}.$
Combining terms again completes the proof. (Recall the second term is negated, and the interpolation bound was computed for $2{\mathrm{GS}}_{3,\varepsilon}(\xi)$.) ∎
Proof of (1.12).
Deducing (1.12) from Corollary 2.7 is identical to (1.11) from the previous subsection. If some ${\boldsymbol{\sigma}}^{1}\in A_{\delta}$ had a small Hessian eigenvalue with eigenvector ${\boldsymbol{v}}$, then setting ${\boldsymbol{\sigma}}^{2},{\boldsymbol{\sigma}}^{3}$ to be $\cos(2\theta(\varepsilon)){\boldsymbol{\sigma}}^{1}\pm\sin(2\theta(\varepsilon)){\boldsymbol{v}}$ similarly to (2.7) would violate the interpolation bound just proved. We omit further details. ∎
2.4 Proof of Corollary 1.7
Here we explain how to deduce transport disorder chaos from Corollary 1.5.
Proof of Corollary 1.7.
The proof essentially follows from [AMS23a, Theorem 5.1], which deduces transport disorder chaos from shattering. Let us summarize the argument. For $\delta\leq\delta_{*}(\xi)$, Corollary 1.5 provides a decomposition $\{{\mathcal{C}}_{-K},\dots,{\mathcal{C}}_{-1},{\mathcal{C}}_{1},\dots,{\mathcal{C}}_{K}\}$ of $A_{\delta}$ into separated clusters, where without loss of generality ${\mathcal{C}}_{-i}=-{\mathcal{C}}_{i}$ are antipodal pairs (since $\xi$ is even). The clusters ${\mathcal{C}}_{i}$ have small diameter $O(\sqrt{N\delta})$ and are pair-wise separated by a much larger distance $\sqrt{N}/C(\xi)$. Further, Lemma 1.12 implies that
$\mu_{\beta}\big{(}\bigcup_{i}{\mathcal{C}}_{i}\big{)}\geq 1-e^{-cN}$ (2.10)
when $\beta\geq\beta(\xi,\delta)$ is very large.
The idea of [AMS23a, Theorem 5.1] is that $\mu_{\beta,\varepsilon_{N}}$ will preserve the clustering (i.e. the property (2.10)), but injects noise into the cluster weights. This is implemented in Proposition 5.4 therein, by proving anti-concentration of the log-ratios $\log\big{(}\mu_{\beta,\varepsilon_{N}}({\mathcal{C}}_{i})/\mu_{\beta,\varepsilon_{N}}({\mathcal{C}}_{j})\big{)}$ for each distinct pair $i\neq j$. The anti-concentration comes from the contribution of ${\boldsymbol{v}}^{\otimes p}$ in $\tilde{H}_{N}$, where $\gamma_{p}>0$ and ${\boldsymbol{v}}$ is chosen such that ±$\pm{\mathcal{C}}_{i},\pm{\mathcal{C}}_{j}$ remain separated when projected along ${\boldsymbol{v}}$. The only input hypothesis that is unavailable in our setting is denoted $\mathsf{S1}$ therein: it is not true that $\mu_{\beta}({\mathcal{C}}_{i})$ is exponentially small for each $i$. However, the proof (in the case of even models) applies with only cosmetic changes as long as ${\mathbb{E}}[\max_{i}\mu_{\beta}({\mathcal{C}}_{i}\cup{\mathcal{C}}_{-i})]\leq 1-C(\xi)^{-1}$, i.e. with uniformly positive probability, the maximum probability of any cluster and its antipodal pair is bounded away from $1$. For this, it is sufficient to show that, if ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\sim\mu_{\beta}$ are i.i.d. Gibbs samples, then
$\liminf_{N\to\infty}\leavevmode\nobreak\ {\mathbb{E}}^{H_{N}}{\mathbb{P}}^{{\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}}[|R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})|\leq 1/2]>0.$ (2.11)
Indeed, $|R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})|\leq 1/2$ implies ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}$ are not in the same cluster, nor in antipodally opposite clusters ${\mathcal{C}}_{i}$ and ${\mathcal{C}}_{-i}$.
To show (2.11) we use the “even generic” hypothesis $\sum_{p\in 2{\mathbb{N}}}\frac{1_{\gamma_{p}\neq 0}}{p}=\infty$. It is standard from [Pan13b, Chapter 3.7], see also [Pan16, Section 4], that under this hypothesis, the law of $|R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})|$ (averaged over the disorder $H_{N}$) has an $N\to\infty$ limit $\zeta_{\beta}$ given by the minimizer in the Parisi formula at inverse temperature $\beta$. The precise definition of $\zeta_{\beta}$ is recalled in Section 3, but the only property we need is the well-known fact that $0\in{\rm supp}(\zeta_{\beta})$ whenever $\xi^{\prime}(0)=0$. In particular, (2.11) holds, so the proof of [AMS23a, Theorem 5.1] applies and yields Corollary 1.7. ∎
3 Alternate Proof of Marginal Stability for Generic Models
Here we give a different proof that full RSB near overlap $1$ implies marginal stability at low temperature. This means we consider the typical behavior of ${\boldsymbol{\sigma}}_{\beta}$ drawn from the Gibbs measure $\mu_{\beta}=\mu_{\beta,H_{N}}$ defined by
${\rm d}\mu_{\beta}({\boldsymbol{\sigma}})=e^{\beta H_{N}({\boldsymbol{\sigma}})}{\rm d}\mu_{0}({\boldsymbol{\sigma}})/Z_{N,\beta}.$
Here $Z_{N,\beta}=\int_{{\mathcal{S}}_{N}}e^{\beta H_{N}({\boldsymbol{\sigma}})}{\rm d}\mu_{0}({\boldsymbol{\sigma}})$ is the partition function relative to uniform measure $\mu_{0}$. The argument is based on ultrametricity of Gibbs measures, which requires $\xi$ to be generic, i.e.
$\sum_{\text{odd}\leavevmode\nobreak\ p}\frac{1_{\gamma_{p}\neq 0}}{p}=\sum_{\text{even}\leavevmode\nobreak\ p}\frac{1_{\gamma_{p}\neq 0}}{p}=\infty.$
(Note the similarity to the “even generic” hypothesis of Corollary 1.7, which would also suffice here.) We prove the following, which is a special case of (1.13) in Corollary 1.4.
Proposition 3.1.
Suppose $\xi$ is generic and not quadratic, and $1$ is not an isolated point in $T$ (recall (1.25)). Then for any $\varepsilon>0$, if ${\boldsymbol{\sigma}}_{\beta}\sim\mu_{\beta}$ for $\beta\geq\beta_{*}(\xi,\varepsilon)$ sufficiently large, and $\delta,c$ are small depending on $(\xi,\varepsilon,\beta)$, and $N$ is sufficiently large, we have with probability $1-e^{-cN}$:
$\|\nabla H_{N}({\boldsymbol{\sigma}}_{\beta})-2\sqrt{\xi^{\prime\prime}(1)}\,{\boldsymbol{\sigma}}_{\beta}\|$ $\leq\varepsilon\sqrt{N},$ (3.1)
$|{\boldsymbol{\lambda}}_{1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}_{\beta})\big{)}|+|{\boldsymbol{\lambda}}_{\lfloor\delta N\rfloor}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}_{\beta})\big{)}|$ $\leq\varepsilon.$ (3.2)
The positive-temperature Parisi formula will enter only by prescribing the possible overlaps of Gibbs samples, giving a more conceptual explanation for this implication. We now recall its statement, which gives the limiting value of the free energy $F_{N,\beta}=\frac{1}{N}\log Z_{N,\beta}$.
Let ${\cal M}$ denote the space of all right-continuous non-decreasing functions $x:[0,1]\to[0,1]$ with $x(\hat{q})=1$ for some $\hat{q}<1$ (which may depend on $x$). Let $\hat{x}(q)=\int_{q}^{1}x(s)\leavevmode\nobreak\ {\rm d}s$ and define the Crisanti–Sommers functional
${\mathcal{P}}(x;\xi)=\frac{1}{2}\left\{\xi^{\prime}(0)\hat{x}(0)+\int_{0}^{1}\xi^{\prime\prime}(q)\hat{x}(q)\leavevmode\nobreak\ {\rm d}q+\int_{0}^{\widehat{q}}\frac{{\rm d}q}{\hat{x}(q)}+\log(1-{\widehat{q}})\right\}.$ (3.3)
Note that $\hat{x}(q)=1-q$ for $q>{\widehat{q}}$, so this functional is independent of ${\widehat{q}}$. The spherical Parisi formula at positive temperature is as follows.
Proposition 3.2 ([Tal06a, Che13, CS17]).
For $\beta\in{\mathbb{R}}_{+}$, the asymptotic free energy satisfies:
$F(\beta)$ $\equiv\operatorname*{p-lim}_{N\to\infty}F_{N,\beta}=\inf_{(\zeta,b)}{\mathcal{P}}_{\beta}(\zeta,b),$
${\mathrm{GS}}$ $=\lim_{\beta\to\infty}F(\beta)/\beta.$
Moreover the minimizers $(\zeta_{\beta},b_{\beta})$ exist and are unique. With $(\zeta,L)$ the zero-temperature minimizers, one has
$\zeta=\lim_{\beta\to\infty}\beta\zeta_{\beta},\quad\quad L=\lim_{\beta\to\infty}\int_{0}^{1}\beta\zeta_{\beta}(s)\leavevmode\nobreak\ {\rm d}s.$
Here $\zeta$ and $\zeta_{\beta}$ are metrized by the vague topology on $[0,1)$ for the corresponding positive measures as in (1.24).
There is a positive-temperature analog of Proposition 1.13 characterizing $(\zeta_{\beta},b_{\beta})$, but this will not be needed. We instead use the following qualitative result on the overlaps between Gibbs samples. Here and below, ${\rm supp}(\zeta_{\beta})\subseteq[0,1)$ is the set of points of increase of $\zeta_{\beta}$.
Corollary 3.3.
Assume the conditions of Proposition 3.1. Then there exists $\varepsilon=\varepsilon_{*}(\xi)>0$ such that for any $q\in(1-\varepsilon_{*},1)$ and $\delta>0$, if $\beta\geq\beta_{*}(\xi,q,\delta)$ is sufficiently large, then ${\rm supp}(\zeta_{\beta})\cap(q-\delta,q+\delta)\neq\emptyset$.
Proof.
As stated in (3.2), the zero-temperature optimizer $\zeta$ is the vague limit as $\beta\to\infty$ of $\beta\zeta_{\beta}$. Hence it suffices to show that $\zeta$ is strictly increasing in a neighborhood of $1$. Proposition 1.13 implies that $T$ contains a non-trivial interval $(1-\varepsilon_{*}(\xi),1)$. On this interval, we must have $g^{\prime\prime}(q)=0$, which easily rearranges to
$\zeta^{\prime}(q)=\big{(}1/\sqrt{\xi^{\prime\prime}(q)}\big{)}^{\prime\prime}.$
The latter is a analytic function on an open complex neighborhood of $[1-\varepsilon_{*},1]$ and is not identically zero since $\xi$ is not quadratic. Thus it has finitely many zeros; this completes the proof. ∎
We consider the event that a large number $K$ of Gibbs samples have all overlaps approximately equal to $q$:
[…]
Here ${\boldsymbol{\sigma}}_{\beta}^{1},\dots,{\boldsymbol{\sigma}}_{\beta}^{K}\stackrel{{\scriptstyle\text{i.i.d.\leavevmode\nobreak\ }}}{{\sim}}\mu_{\beta,H_{N}}$ are always i.i.d. samples from $\mu_{\beta}$. Our alternate proof of Proposition 3.1 relies on the fact that $E_{K,q,\delta}$ has uniformly positive probability when $q\in{\rm supp}(\zeta_{\beta})$, even conditional on typical $(H_{N},{\boldsymbol{\sigma}}_{\beta}^{1})$. This property follows from the Ghirlanda-Guerra identities, which hold when $\xi$ is generic. Namely as explained in [Pan13b, Chapter 3.7], when $\xi$ is generic the limiting overlap arrays of i.i.d. Gibbs samples converge to a limiting random overlap structure as $N\to\infty$. It is easy to see that this limiting structure satisfies the next proposition.
Proposition 3.4.
If $\xi$ is generic, then for all $q\in{\rm supp}(\zeta_{\beta})$ and any $\delta>0$,
$\lim_{\eta\to 0}\lim_{N\to\infty}{\mathbb{P}}\left[{\mathbb{P}}[E_{K,q,\delta}\leavevmode\nobreak\ |\leavevmode\nobreak\ (H_{N},{\boldsymbol{\sigma}}_{\beta}^{1})]>\eta\right]=1.$
3.1 Preparatory Lemmas
The next proposition shows that low dimensional spaces do not contain a large number of points with all equal distances. Though elementary, it is key to our argument.
Proposition 3.5.
For any $d>0$ there exists $\varepsilon>0$ such that no $d+2$ points $x_{1},\dots,x_{d+2}\in{\mathbb{R}}^{d}$ satisfy
[…]
for any $a>0$.
Proof.
Without loss of generality set $a=1$. In the proof below, implicit constants in $O(\cdot)$ may depend on $d$. Suppose such points exist and let $z=\frac{1}{d+2}\sum_{i=1}^{d+2}x_{i}$ and $y_{i}=x_{i}-z$. Note that for each $j,k\neq i$:
$\langle x_{i}-x_{j},x_{i}-x_{k}\rangle=\frac{\|x_{i}-x_{j}\|^{2}+\|x_{i}-x_{k}\|^{2}-\|x_{j}-x_{k}\|^{2}}{2}=\frac{1}{2}\pm O(\varepsilon).$
We compute
$\langle y_{i},y_{i}\rangle=(d+2)^{-2}\sum_{k,\ell=1}^{d+2}\langle x_{i}-x_{k},x_{i}-x_{\ell}\rangle\\ =\left(\frac{d+1}{d+2}\right)^{2}\pm O(\varepsilon).$
Similarly $|\langle x_{i}-x_{k},x_{j}-x_{\ell}\rangle|\leq O(\varepsilon)$ when all four indices are distinct, so for $i\neq j$:
$\langle y_{i},y_{j}\rangle=(d+2)^{-2}\sum_{k,\ell=1}^{d+2}\langle x_{i}-x_{k},x_{j}-x_{\ell}\rangle=-\frac{d+1}{(d+2)^{2}}\pm O(\varepsilon).$
Hence the $(d+2)\times(d+2)$ matrix $M$ with entries $M_{i,j}=\langle y_{i},y_{j}\rangle$ is entrywise within $O(\varepsilon)$ of
$\widetilde{M}_{i,j}=\begin{cases}\left(\frac{d+1}{d+2}\right)^{2},\quad i=j,\\ -\frac{d+1}{(d+2)^{2}},\quad i\neq j.\end{cases}$
Diagonal dominance implies ${\rm rank}(\widetilde{M})=d+1$, hence ${\rm rank}(M)\geq d+1$ for $\varepsilon$ sufficiently small. However by construction ${\rm rank}(M)\leq d$ since $y_{1},\dots,y_{d+2}\in{\mathbb{R}}^{d}$. This is a contradiction and completes the proof. ∎
3.2 Proof of Proposition 3.1
In light of Lemma 1.11, it will suffice to prove (3.2) with $\lfloor\delta N\rfloor$ replaced by a large constant $K$. Thus, for sake of contradiction we fix $K,C>0$ such that
$\limsup_{\beta\to\infty}\limsup_{N\to\infty}{\mathbb{P}}[{\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))\leq-C]>0.$ (3.5)
We next choose several more constants: $\varepsilon\leq\varepsilon_{*}(\xi,K,C)$ is taken sufficiently small such that $q=1-\varepsilon\in{\rm supp}(\zeta)$, and we set $\lambda=\varepsilon^{0.55}$. Then we send $\beta\to\infty$, inducing a choice of $\eta\to 0$ so that Corollary 1.10 holds, and $\delta\to 0$ so Corollary 3.3 holds. Thus $\beta$ (resp. $\eta,\delta)$ is sufficiently large (resp. small) depending on $(\xi,K,C,\varepsilon)$.
Given ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$, let ${\mathcal{T}}({\boldsymbol{\sigma}})$ be the tangent space to ${\mathcal{S}}_{N}$, viewed as a codimension $1$ linear subspace of ${\mathbb{R}}^{N}$. Let $U_{K}=U_{K}({\boldsymbol{\sigma}})\subseteq{\mathcal{T}}({\boldsymbol{\sigma}})$ denote the span of the top $K$ eigenvectors of $\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}})$, and let $U_{K}^{\perp}({\boldsymbol{\sigma}})\subseteq{\mathcal{T}}({\boldsymbol{\sigma}})$ be its orthogonal complement in ${\mathcal{T}}({\boldsymbol{\sigma}})$. For any ${\boldsymbol{v}}\in{\mathbb{R}}^{N}$, there is a unique decomposition
${\boldsymbol{v}}={\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}+R({\boldsymbol{\sigma}},v){\boldsymbol{\sigma}}$ (3.6)
with $({\boldsymbol{v}}_{\parallel},{\boldsymbol{v}}_{\perp})\in U_{K}\times U_{K}^{\perp}$. Note that if ${\boldsymbol{v}}={\boldsymbol{\sigma}}^{\prime}-{\boldsymbol{\sigma}}$ then ${\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}$ is proportional to the derivative $\gamma^{\prime}(0)$ for $\gamma$ a geodesic path from ${\boldsymbol{\sigma}}$ to ${\boldsymbol{\sigma}}^{\prime}$. We also set $U_{K,\lambda}$ to be the $\lambda\sqrt{N}$-neighborhood of $U_{K}({\boldsymbol{\sigma}})+{\boldsymbol{\sigma}}$ in ${\mathbb{R}}^{N}$ (note that $U_{K}\subseteq{\mathcal{T}}({\boldsymbol{\sigma}})\subseteq{\mathbb{R}}^{N}$ can be naturally viewed as a subset of ${\mathbb{R}}^{N}$).
Proposition 3.6.
For ${\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime}\in{\mathcal{S}}_{N}$, let ${\boldsymbol{v}}={\boldsymbol{\sigma}}^{\prime}-{\boldsymbol{\sigma}}$ and define the decomposition (3.6) based on $U_{K}({\boldsymbol{\sigma}})$. Then uniformly over $H_{N}\in K_{N}$:
$H_{N}({\boldsymbol{\sigma}}^{\prime})=H_{N}({\boldsymbol{\sigma}})+\langle\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}),{\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}\rangle+\langle\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}),({\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp})^{\otimes 2}\rangle+N^{-1/2}\cdot O(\|{\boldsymbol{\sigma}}^{\prime}-{\boldsymbol{\sigma}}\|^{3}).$
Proof.
This amounts to a second order Taylor expansion of $H_{N}$ along the geodesic path from ${\boldsymbol{\sigma}}$ to ${\boldsymbol{\sigma}}^{\prime}$. The required estimate on the third derivative holds since $H_{N}\in K_{N}$. ∎
We now show the Gibbs mass near ${\boldsymbol{\sigma}}_{\beta}$ essentially lives within the set $U_{K,\lambda}({\boldsymbol{\sigma}}_{\beta})$.
Lemma 3.7.
For constants chosen as above, suppose $H_{N}\in K_{N}$ and that ${\boldsymbol{\sigma}}\in{\mathcal{S}}_{N}$ satisfies:
$\|\nabla H_{N}({\boldsymbol{\sigma}})\|\leq\eta\sqrt{N};\quad\quad{\boldsymbol{\lambda}}_{1}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}))\leq\eta;\quad\quad{\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}))\leq-C.$
Then it follows that
$\mu_{\beta}\big{(}B_{2\sqrt{\varepsilon N}}({\boldsymbol{\sigma}})\backslash U_{K,\lambda}({\boldsymbol{\sigma}})\big{)}\leq e^{-cN}.$
Proof.
Given ${\boldsymbol{\sigma}}^{\prime}\in B_{2\sqrt{\varepsilon N}}({\boldsymbol{\sigma}})$, let ${\boldsymbol{v}}={\boldsymbol{\sigma}}^{\prime}-{\boldsymbol{\sigma}}$ and let ${\boldsymbol{w}}={\boldsymbol{v}}_{\parallel}+{\boldsymbol{v}}_{\perp}$ be the associated tangent vector to ${\mathcal{S}}_{N}$ at ${\boldsymbol{\sigma}}$ with $\|{\boldsymbol{w}}\|\leq 2\sqrt{\varepsilon N}$. From the way we chose constants, $\lambda^{2}$ is larger than $\eta\varepsilon^{1/2}+\varepsilon^{3/2}$ by a super-constant factor. Hence Proposition 3.6 yields
$H_{N}({\boldsymbol{\sigma}}^{\prime})$ $\leq H_{N}({\boldsymbol{\sigma}})+\eta(\|{\boldsymbol{w}}\|\sqrt{N}+\|{\boldsymbol{w}}\|^{2})-C\|{\boldsymbol{v}}^{\perp}\|^{2}/2+O(\varepsilon^{3/2}N)$
$\leq H_{N}({\boldsymbol{\sigma}})+O(\eta\varepsilon^{1/2}+\varepsilon^{3/2})N-C\|{\boldsymbol{v}}^{\perp}\|^{2}/2\leq H_{N}({\boldsymbol{\sigma}})-C\lambda^{2}N/3.$
Lemma 1.12 now completes the proof. ∎
Lemma 3.8.
With parameters as above (in particular $\delta$ small depending on $\varepsilon$), suppose ${\boldsymbol{\sigma}}^{1},\dots,{\boldsymbol{\sigma}}^{K+3}\in{\mathcal{S}}_{N}$ satisfy:
[…]
Then with decomposition (3.6) defined based on $U_{K}({\boldsymbol{\sigma}}^{1})$, and with ${\boldsymbol{v}}^{j}={\boldsymbol{\sigma}}^{j}-{\boldsymbol{\sigma}}^{1}$ for $32\leq j\leq K+$ 3,
$\max_{2\leq j\leq K+3}\|{\boldsymbol{v}}^{j}_{\perp}\|>\lambda\sqrt{N}.$
Proof.
Suppose not. Then the vectors ${\boldsymbol{v}}^{j}_{\parallel}$ for $32\leq j\leq K+$ 3 lie in a $K$-dimensional subspace and for $i\neq j$:
$\|{\boldsymbol{v}}^{i}_{\parallel}-{\boldsymbol{v}}^{j}_{\parallel}\|$ $=\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{j}\|\pm O(\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{i}_{\parallel}\|+\|{\boldsymbol{v}}^{j}-{\boldsymbol{v}}^{j}_{\parallel}\|)=\|{\boldsymbol{\sigma}}^{i}-{\boldsymbol{\sigma}}^{j}\|\pm O(\|{\boldsymbol{v}}^{i}-{\boldsymbol{v}}^{i}_{\parallel}\|+\|{\boldsymbol{v}}^{j}-{\boldsymbol{v}}^{j}_{\parallel}\|)$
$=\Big{(}\sqrt{2-2(1-\varepsilon)^{2}}\pm O(\delta\varepsilon^{-1/2}+\lambda)\Big{)}\sqrt{N}.$
This is impossible by Proposition 3.5 since $\sqrt{2-(2-\varepsilon)^{2}}\asymp\sqrt{\varepsilon}$ and $\delta\varepsilon^{-1/2}+\lambda\leq o(\sqrt{\varepsilon})$, completing the proof. ∎
Proof of Proposition 3.1.
First, Corollary 1.10 implies $\|\nabla_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta})\|\leq\varepsilon\sqrt{N}/2$. It suffices to establish (3.2) with $\lfloor\delta N\rfloor$ replaced by $K$, as Lemma 1.11 shows these are equivalent and also then yields (3.1). Combining Lemma 1.11 with (1.22) gives the upper bound for ${\boldsymbol{\lambda}}_{1}\big{(}\nabla_{\mathrm{sp}}^{2}H_{N}({\boldsymbol{\sigma}}_{\beta})\big{)}$.
The main part of the proof is the lower bound on ${\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))$. Suppose for sake of contradiction that (3.5) holds for some $K,C>0$, and let $S_{{\rm eigen}}$ denote the event that ${\boldsymbol{\lambda}}_{K}(\nabla^{2}_{\mathrm{sp}}H_{N}({\boldsymbol{\sigma}}_{\beta}))\leq-C$. Choose small $\varepsilon$ depending on $(K,C)$ such that $1-\varepsilon\in{\rm supp}(\zeta)$. Let $\eta\ll\delta$ be small depending on $(\xi,K,C,\varepsilon,\beta)$, and define for i.i.d. Gibbs samples $({\boldsymbol{\sigma}}_{\beta}^{1},\dots,{\boldsymbol{\sigma}}_{\beta}^{K+3})$ the event
$S_{{\rm generic}}=\big{\{}{\mathbb{P}}[E_{K+3,1-\varepsilon,\delta}\leavevmode\nobreak\ |\leavevmode\nobreak\ (H_{N},{\boldsymbol{\sigma}}_{\beta}^{1})]>(K+2)\eta\big{\}}.$
By Proposition 3.4, for $N\geq N_{0}(\xi,K,C,\varepsilon,\beta,\delta,\eta)$ sufficiently large we have
${\mathbb{P}}\left[S_{{\rm generic}}\right]>1-\frac{\varepsilon}{2}.$
The event $E_{K+3,1-\varepsilon,\delta}$ trivially implies ${\boldsymbol{\sigma}}_{\beta}^{i}\in B_{2\sqrt{\varepsilon N}}({\boldsymbol{\sigma}}_{\beta}^{1})$ for all $32\leq i\leq K+$ 3 because $\delta$ is small depending on $\varepsilon$. Explicitly, if $R({\boldsymbol{\sigma}},{\boldsymbol{\sigma}}^{\prime})\geq 1-\varepsilon-\delta$ then $\|{\boldsymbol{\sigma}}-{\boldsymbol{\sigma}}^{\prime}\|=\sqrt{2(\varepsilon+\delta)N}\leq 2\sqrt{\varepsilon N}$. Let $S_{{\rm bounded}}$ be the event $H_{N}\in K_{N}$, which has probability $1-e^{-cN}$ by Proposition 1.9. We claim the three events $S_{*}$ cannot all hold, i.e. deterministically,
$S_{{\rm generic}}\cap S_{{\rm bounded}}\cap S_{{\rm eigen}}=\emptyset.$ (3.7)
Indeed, assume that $S_{{\rm generic}}$ and $S_{{\rm bounded}}$ hold. Then using Lemma 3.8 in the first step and the definition of $S_{{\rm generic}}$,
$\mu_{\beta}\big{(}B_{2\sqrt{\varepsilon N}}({\boldsymbol{\sigma}}_{\beta}^{1})\backslash U_{K,\lambda}({\boldsymbol{\sigma}}^{1})\big{)}\geq{\mathbb{P}}[E_{K+3,1-\varepsilon,\delta}\leavevmode\nobreak\ |\leavevmode\nobreak\ (H_{N},{\boldsymbol{\sigma}}_{\beta}^{1})]/(K+2)\geq\eta.$
In light of Lemma 3.7 and the assumption $S_{{\rm bounded}}$, we find that $S_{{\rm eigen}}$ indeed cannot hold. Finally it follows from (3.7) that ${\mathbb{P}}[S_{{\rm eigen}}]\leq\frac{\varepsilon}{2}+e^{-cN}\leq\varepsilon$ which concludes the proof. ∎
Acknowledgements
Thanks to Brice Huang for helpful discussions, and for comments on a previous version of the paper. We also thank Andrea Montanari and Pierfrancesco Urbani for enlightening conversations.